### v1

Abstract
Various tasks in scientific computing can be modeled as an optimization problem on the indefinite Stiefel manifold. We address this using the Riemannian approach, which basically consists of equipping the feasible set with a Riemannian metric, preparing geometric tools such as orthogonal projections, formulae for Riemannian gradient, retraction and then extending an unconstrained optimization algorithm on the Euclidean space to the established manifold. The choice for the metric undoubtedly has a great influence on the method. In the previous work [D.V. Tiep and N.T. Son, A Riemannian gradient descent method for optimization on the indefinite Stiefel manifold, arXiv:2410.22068v2[math.OC]], a tractable metric, which is indeed a family of Riemannian metrics defined by a symmetric positive-definite matrix depending on the contact point, has been used. In general, it requires solving a Lyapunov matrix equation every time when the gradient of the cost function is needed, which might significantly contribute to the computational cost. To address this issue, we propose a new Riemannian metric for the indefinite Stiefel manifold. Furthermore, we construct the associated geometric structure, including a so-called quasi-geodesic and propose a retraction based on this curve. We then numerically verify the performance of the Riemannian gradient descent method associated with the new geometry and compare it with the previous work.
1 Introduction
Optimization of smooth functions of matrix variable $X$ subject to the constraint of the form $\mathcal{M}=\{X\ :\ X^{T}AX=J\}$, where $A$ is nonsingular and symmetric and $J$ is symmetric, involutory, i.e., satisfying $J^{2}=I$, is an established line of researches due to its wide range of applications. Various problems such as dimensionality reduction of data via principal component analysis, canonical correlation analysis, finding locality preserving projection, and computation of the generalized eigenvalues of a positive-definite matrix pencil can be formulated as an optimization problem with such a constraint, see, e.g., HardSST04 ; HeN03 ; Hote92 ; KovaV95 ; SameW82 ; TiepSon2025 ; YgerBGR12 .
Additional conditions can be imposed on the matrices $A$ and $J$ upon the applications considered making the feasible set a differentiable manifold. When $A=I_{n}$ and $J=I_{k},k\leq n$, the $n\times n$ and $k\times k$ identity matrices, respectively, $\mathcal{M}$ is termed the (orthogonal) Stiefel manifold EdelAS98
$\mathrm{St}(k,n):=\{X\in{\mathbb{R}^{n\times k}}\ :\ X^{T}X=I_{k}\}.$
Alternatively, if $A$ is an $n\times n$ real symmetric positive-definite (spd) matrix and $J=I_{k}$, the set is called the generalized Stiefel manifold
$\mathrm{St}_{A}(k,n):=\{X\in{\mathbb{R}^{n\times k}}\ :\ X^{T}AX=I_{k}\},$
see, e.g., ShusA2023 . Although $\mathrm{St}(k,n)$ is a special case of $\mathrm{St}_{A}(k,n)$, one can see that $\mathrm{St}_{A}(k,n)$ can be turned into a Stiefel manifold by the mapping $X\mapsto R^{-1}Y$ where $A=R^{T}R$ is the Cholesky factorization of $A$. A more general one in which $A$ is nonsingular, symmetric, and $J$ is symmetric, involutory, which covers the orthogonal and generalized Stiefel manifolds as special cases, is the indefinite Stiefel manifold
$\mathrm{iSt}_{A,J}(k,n):=\{X\in{\mathbb{R}^{n\times k}}\ :\ X^{T}AX=J\}.$ (1)
This manifold was recently introduced in TiepSon2025 . Exploiting the rich geometry of the feasible set, a popular approach, known as Riemannian optimization, is to extend well-known methods for unconstrained optimization in the Euclidean space to the case of these Riemannian manifolds once necessary geometric tools such as a Riemannian metric, orthogonal projections, a retraction are constructed. One is referred to classical works AdleDMMS02 ; EdelAS98 , monographs AbsiMS08 ; Boumal23 ; Sato2021 ; Udri1994 , and articles manopt ; MishS2016 ; SatoA19 ; ShusA2023 ; TiepSon2025 ; WangDPY2024 ; YgerBGR12 , to name a few among many others dedicated to optimization methods on the orthogonal, generalized, and indefinite Stiefel manifolds.
Equipping a Riemannian metric for a differentiable manifold is an essential step in Riemannian optimization. Recall that a Riemannian metric is an inner product defined on each tangent space of the manifold depending smoothly on the contact point. It not only helps quantify objects such as the length of vectors, the measure of angles but also has a great effect on the behavior of numerical methods. Specifically, it is a means of defining the Riemannian gradient of a cost function, a crucial material in most of optimization methods.
There are different ways to equip a manifold with a Riemannian metric depending on how the manifold is viewed. For some left/right-invariant Riemannian metrics, pseudo-Riemannian metrics equipped for the Stiefel manifold as a quotient manifold of two orthogonal groups, one is referred to (EdelAS98, , Subsect. 2.3), JurdML2020 ; NishA2005 ; cf. BendZ21 for the symplectic Stiefel manifold. Here, we follow (TiepSon2025, , Prop. 2.2) in which $\mathrm{iSt}_{A,J}(k,n)$ is shown to be an embedded submanifold of $\mathbb{R}^{n\times k}$} of dimension $nk-\frac{1}{2}k(k+1)$. Then, a Riemannian metric is defined via an $n\times n$ spd $X$-smoothly depending matrix $\mathbf{M}_{X}$ in which the inner product of $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}_{A,J}(k,n)$ is determined as
$\mathrm{g}_{\mathbf{M}_{X}}(Z_{1},Z_{2}):=\mathrm{tr}\left(Z_{1}^{T}\mathbf{M}_{X}Z_{2}\right).$
This is not the most general way of defining a metric on a submanifold of $n\times k$ structured matrices but it covers most cases in practice, e.g., the standard Euclidean metric and the canonical metric EdelAS98 . Using this formulation, the Riemannian gradient of smooth cost functions via orthogonally projecting onto the tangent space requires solving a Lyapunov matrix equation (TiepSon2025, , Prop. 3.3), which can be computationally expensive when $k$ is not small. It is surprising that in the case of Stiefel manifold, the Riemannian gradient with respect to the Euclidean and canonical metrics have closed-form expressions, see (EdelAS98, , Sect. 2) and (AbsiMS08, , Subsect. 3.6). Moreover, in the generalized Stiefel manifold $\mathrm{St}_{A}(k,n)$ when $A$ is chosen to be the representing matrix $\mathbf{M}_{X}$, a closed-form Riemannian gradient is available (WangDPY2024, , Sect. 4). Meanwhile, for the symplectic Stiefel manifold, a relatively similar manifold to $\mathrm{St}_{A}(k,n),$ computing the Riemannian gradient needs solving a Lyapunov matrix equation under the Euclidean metric and a general tractable metric and does not when the canonical-like metric is used GSAS21a ; GSAS21 ; GSS2024_2 . On the one hand, these facts raise a hope that one can design a Riemannian metric, most probably the canonical(-like) metric as in EdelAS98 ; GSAS21 , for the indefinite Stiefel manifold such that solving the Lyapunov matrix equation is not compulsory in order to compute the Riemannian gradient. Note, however, that the motivation for using the canonical(-like) metric in EdelAS98 ; GSAS21 is to balance the contribution of the two components of tangent vectors in the metric formulation instead of obviating the necessity of solving the Lyapunov matrix equation as that for the metric proposed in this work. On the other hand, they inspire several questions: (1) How are we on track for not having to solve the Lyapunov matrix equations to compute the Riemannian gradient, (2) How to explain the different situations for orthogonal, generalized, and symplectic Stiefel manifolds mentioned above, and (3) How does an optimization method such as Riemannian gradient descent behave using the geometric quantities under the proposed metric?
Contribution
It is the main motivation source of the present paper to answer these questions. Namely, by converting the Lyapunov matrix equation into a linear equation using the Kronecker product and vectorization operator and by asking that the resulting equation is solving-free, we first achieve a general one-parameter family of metrics that fulfill our expectation; the canonical(-like) metric in EdelAS98 ; GSAS21 certainly belongs to this class. Second, we explain in detail why in the case of the orthogonal Stiefel manifold, the Lyapunov matrix equation does not appear during the computation of the Riemannian gradient by examining the formulation of the orthogonal projection onto the tangent space under the tractable metric. Third, we go further by characterizing the normal space to the manifold, deriving closed-form expressions for the orthogonal projections onto the tangent and normal spaces, and the Riemannian gradient of the cost function. Moreover, we construct a so-called quasi-geodesic curve associated with the obtained metric which in turn allows us to propose a retraction. Finally, these tools are invoked in a Riemannian gradient descent scheme for running several numerical examples.
Outline
After recalling some main facts about the indefinite Stiefel manifold in Section 2, we derive the generalized canonical metric that yields closed-form expressions for the orthogonal projections and the Riemannian gradient in Section 3. An explanation for the aforementioned situations in orthogonal, generalized, and symplectic Stiefel manifolds is also given. In Section 4, we construct a quasi-geodesic which is used to propose a retraction in Section 5. In Section 6, using a Riemannian gradient descent algorithm, we verify the theoretical findings via several examples. Finally, a conclusion is given in Section 7.
Notation
We finish this section by introducing the notations. We use $\Omega^{T},\mathrm{tr}(\Omega),$ $\|\Omega\|_{F},$ and $\|\Omega\|$ to signify the transpose, trace, Frobenius norm, and spectral norm of a matrix $\Omega,$ respectively. When $\Omega$ is a square matrix, its skew-symmetric and symmetric parts are denoted by $\mathrm{skew}(\Omega)\coloneqq\frac{1}{2}(\Omega-\Omega^{T})$ and $\mathrm{sym}(\Omega)\coloneqq\frac{1}{2}(\Omega+\Omega^{T}),$ respectively. The set of skew-symmetric, the set of symmetric, and the set of symmetric positive-definite matrices of size $p\times p$ are respectively denoted by $\mathrm{Skew}(p)$, $\mathrm{Sym}(p)$, and $\mathrm{SPD}(p)$. We write $\mathrm{diag}(\cdot,\ldots,\cdot)$ to mean the (block) diagonal matrix. The matrix exponential is represented by either $\mathrm{e}^{\cdot}$ or $\mathrm{exp}(\cdot).$ In a linear space, $\langle\cdot,\cdot\rangle$ denotes the standard Euclidean inner product; specifically, $\langle P,Q\rangle=\mathrm{tr}(P^{T}Q)$ for $P,Q\in{{\mathbb{R}^{n\times k}}}$. The first-order derivative with respect to $t$ of a smooth mapping $G(t)$ is represented as $\frac{\mathrm{d}}{\mathrm{d}t}G$ or $\dot{G}(t),$ and $\ddot{G}(t)$ means its second-order derivative. For a map $F:\mathcal{V}_{1}\longrightarrow\mathcal{V}_{2}$ between two normed vector spaces $\mathcal{V}_{1}$ and $\mathcal{V}_{2},$ let $o(\cdot)$ denote the little-$o$ notation, the Fréchet derivative of $F$ at $X\in\mathcal{V}_{1}$ is a linear operator $\textsc{D}F(X):\mathcal{V}_{1}\longrightarrow\mathcal{V}_{2}$ fulfilling that $F(X+Z)=F(X)+\textsc{D}F(X)[Z]+o(\|Z\|),$ for any $Z\in\mathcal{V}_{1}$.
2 Preliminaries
In this section, we recall basic facts about the indefinite Stiefel manifold $\mathrm{iSt}_{A,J}(k,n)$ which can also be shortened as $\mathrm{iSt}_{A,J}$ or $\mathrm{iSt}(k,n)$ if the dropped information is either inessential or clear from the context. For their derivations, the reader is referred to TiepSon2025 .
Given matrix $A$ as introduced in the beginning, let $\mathrm{i}_{+}(A)$ and $\mathrm{i}_{-}(A)$ respectively denote the number of positive and negative eigenvalues of $A$. It was stated in (TiepSon2025, , Lem. 2.1) that a necessary and sufficient condition for the nonemptiness of the feasible set $\mathrm{iSt}_{A,J}$ is that
[…]
We always assume that (2) holds when working with this manifold.
Then, it was shown in (TiepSon2025, , Prop. 2.2) that the set $\mathrm{iSt}_{A,J}$ defined in (1) is a closed, embedded submanifold of $\mathbb{R}^{n\times k}$} of dimension $nk-\frac{1}{2}k(k+1)$. It was also noted that the indefinite Stiefel manifold is generally unbounded.
For $X\in\mathrm{iSt}_{A,J}(k,n)$, let us denote by $X_{\perp}\in\mathbb{R}^{n\times(n-k)}$ a full-rank matrix satisfying $X^{T}X_{\perp}=0\in\mathbb{R}^{k\times(n-k)}$ and set $E\coloneqq[X\;\;A^{-1}X_{\perp}]\in{\mathbb{R}^{n\times n}}.$ The following statements hold due to (TiepSon2025, , Lem. 2.3):
  • i)
    $E$ is nonsingular and
    $E^{-1}=\begin{bmatrix}JX^{T}A\\ (X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}\end{bmatrix}$. (3)
  • ii)
    For any $Z\in{\mathbb{R}^{n\times k}},$ there are $W\in{\mathbb{R}^{k\times k}}$ and $K\in\mathbb{R}^{(n-k)\times k}$ such that
    $Z=XW+A^{-1}X_{\perp}K,$ (4)
    where $W=JX^{T}AZ,K=(X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}Z,$ and $A^{-1}X_{\perp}K$ is independent of $X_{\perp}.$
  • iii)
    There holds
    $I_{n}=XJX^{T}A+A^{-1}X_{\perp}(X_{\perp}^{\top}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}.$ (5)
Next, recall that the tangent space at $X\in\mathrm{iSt}(k,n)$ is the set
$T_{X}\mathrm{iSt}(k,n)=\left\{\dot{\varphi}(0)\;:\;\varphi:(a,b)\ni 0\longrightarrow\mathrm{iSt}(k,n)\text{ is smooth and }\varphi(0)=X\right\}.$
In (TiepSon2025, , Prop. 2.4), it has been shown that
$T_{X}\mathrm{iSt}(k,n)$ $=\left\{Z\in{\mathbb{R}^{n\times k}}:Z^{T}AX+X^{T}AZ=0\right\}$
$=\left\{XW+A^{-1}X_{\perp}K:JW\in\mathrm{Skew}(k),K\in\mathbb{R}^{(n-k)\times k}\right\}.$ (6)
In (TiepSon2025, , Sect. 3), a so-called tractable metric $\mathrm{g}_{\mathbf{M}_{X}}$ defined via an $X$-smoothly dependent matrix $\mathbf{M}_{X}\in\mathrm{SPD}(n)$ was proposed. Namely, for $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}(k,n),$ $\langle Z_{1},Z_{2}\rangle_{\mathrm{g}_{\mathbf{M}_{X}}}\coloneq\mathrm{tr}\left(Z_{1}^{T}\mathbf{M}_{X}Z_{2}\right)$. With this metric, for any $X\in\mathrm{iSt}(k,n)$, the ambient space $\mathbb{R}^{n\times k}$} is decomposed as the direct sum of the tangent space and the normal space,
${\mathbb{R}^{n\times k}}=T_{X}\mathrm{iSt}(k,n)\overset{\perp_{\mathrm{g}_{\mathbf{M}_{X}}}}{\bigoplus}T_{X}\mathrm{iSt}(k,n)^{\perp}_{\mathrm{g}_{\mathbf{M}_{X}}}.$
This feature motivates definitions of the orthogonal projections onto the tangent space and onto the normal space which associate each $Y\in{\mathbb{R}^{n\times k}}$ with a unique $\mathscr{P}_{X}(Y)\in T_{X}\mathrm{iSt}(k,n)$ and a unique $\mathscr{P}_{X}^{\perp}(Y)\in T_{X}\mathrm{iSt}(k,n)^{\perp}_{\mathrm{g}_{\mathbf{M}_{X}}}$, respectively, such that
$Y=\mathscr{P}_{X}(Y)+\mathscr{P}_{X}^{\perp}(Y).$ (7)
The computation of $\mathscr{P}_{X}(Y)$ is essential in formulating the Riemannian gradient. According to (TiepSon2025, , Prop. 3.2), $\mathscr{P}_{X}(Y)=Y-\mathbf{M}_{X}^{-1}AXU_{X,Y}$ where $U_{X,Y}$ is the solution to the Lyapunov matrix equation
$2\mathrm{sym}(X^{T}AY)=X^{T}A\mathbf{M}_{X}^{-1}AXU_{X,Y}+U_{X,Y}X^{T}A\mathbf{M}_{X}^{-1}AX.$ (8)
Based on this formula, the Riemannian gradient was given in (TiepSon2025, , Prop. 3.3) by
$\mathrm{grad}_{\mathbf{M}_{X}}f(X)=\mathbf{M}_{X}^{-1}\nabla\bar{f}(X)-\mathbf{M}_{X}^{-1}AXU_{\bar{f}},$ (9)
where $\bar{f}$ is some smooth extension of $f$ on the ambient space $\mathbb{R}^{n\times k}$}, $U_{\bar{f}}$ is the unique solution to the Lyapunov matrix equation with unknown $U$:
$(X^{T}A\mathbf{M}_{X}^{-1}AX)U+U(X^{T}A\mathbf{M}_{X}^{-1}AX)=2\mathrm{sym}\left[X^{T}A\mathbf{M}_{X}^{-1}\nabla\bar{f}(X)\right].$ (10)
Ignoring the computational cost for $\mathbf{M}_{X}^{-1}(\cdot)$ by, e.g., considering the Euclidean metric, one can show that the total cost for computing the Riemannian gradient $\mathrm{grad}_{\mathbf{M}_{X}}f(X)$ in (9) is about $2n^{2}k+6nk^{2}-nk-2k^{2}+20k^{3}$ flops. The last term in this expression is contributed by solving the equation (10) with an spd dense matrix coefficient using the Bartels-Stewart algorithm BartS72 . When $k$ is large, for instance, the same as $n$, it dominates the total cost. Therefore, avoiding solving this equation in such a case most probably helps accelerate the optimization process.
3 A generalized canonical metric on $\mathrm{iSt}(k,n)$
As introduced, we will determine a form of $\mathbf{M}_{X}$ to eliminate the need of solving the Lyapunov equation (10). To this end, we simplify the coefficient $X^{T}A\mathbf{M}_{X}^{-1}AX$ of the equation. The identity $E^{T}AX=\left[\begin{smallmatrix}J\\ 0\end{smallmatrix}\right]$ suggests considering $\mathbf{M}_{X}^{-1}\in\mathrm{SPD}(n)$ as
$\mathbf{M}_{X}^{-1}=E\begin{bmatrix}\Gamma_{1}&\Gamma_{2}\\ \Gamma_{2}^{T}&\Gamma_{3}\end{bmatrix}E^{T},\ \Gamma_{1}\in\mathrm{SPD}(k),\;\Gamma_{2}\in\mathbb{R}^{k\times(n-k)},\;\Gamma_{3}\in\mathrm{SPD}(n-k).$
This results in $X^{T}A\mathbf{M}_{X}^{-1}AX=J\Gamma_{1}J.$ Consequently, the equation (8) reduces to
$2\mathrm{sym}(X^{T}AY)=(J\Gamma_{1}J)U_{X,Y}+U_{X,Y}(J\Gamma_{1}J).$ (11)
For a greater detail, let us convert this matrix equation into a linear equation using vectorization operator and Kronecker product HornJ91 . Namely, for a matrix $M=(m_{ij})\in{\mathbb{R}^{n\times k}}$, $\mathrm{vec}(M)\in\mathbb{R}^{nk}$ is generated by stacking the columns of $M$ to make a column vector and the Kronecker product of $M$ and $N\in\mathbb{R}^{p\times q}$, denoted by $M\otimes N\in\mathbb{R}^{np\times kq}$, is the matrix consisting of $nk$ block matrices $m_{ij}N$ for $i=1,\ldots,n,j=1,\ldots,k$. In view of (HornJ91, , Lem. 4.3.1), (11) is equivalent to
$(I_{k}\otimes(J\Gamma_{1}J)+(J\Gamma_{1}J)\otimes I_{k})\mathrm{vec}(U_{X,Y})=\mathrm{vec}(2\mathrm{sym}(X^{T}AY)).$
We say that a square system of linear equations is solving-free if its coefficient is a scalar matrix, i.e., the product of a (nonzero) scalar and the identity matrix. Thus, its solution is merely the product of the reciprocal of the scalar and the right-hand side. Accordingly, a sufficient condition for obviating the need of solving the Lyapunov matrix equation (8) is that $I_{k}\otimes(J\Gamma_{1}J)+(J\Gamma_{1}J)\otimes I_{k}$ has the form $2\rho I_{k^{2}}$, where $\rho$ is a nonzero constant parameter. It follows that $J\Gamma_{1}J=\rho I_{k}$ which, together with the condition $J^{2}=I_{k}$, yields $\Gamma_{1}=\rho I_{k}$. For the symmetric positive-definiteness of $\mathbf{M}_{X}^{-1}$, we set $\rho>0$. One can verify that in this case, the solution to (11) is $U_{X,Y}=\frac{1}{\rho}\mathrm{sym}(X^{T}AY).$ Furthermore, we set $\Gamma_{2}=0$ for the ease of argument. Consequently, $\mathbf{M}_{X}^{-1}$ takes the form
$\mathbf{M}_{X}^{-1}=E\begin{bmatrix}\rho I_{k}&0\\ 0&\Gamma_{3}\end{bmatrix}E^{T}=\rho XX^{T}+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}.$ (12)
Moreover, taking (3) into account, we also have
$\mathbf{M}_{X}$ $=\begin{bmatrix}JX^{T}A\\ (X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}\end{bmatrix}^{T}\begin{bmatrix}\frac{1}{\rho}I_{k}&0\\ 0&\Gamma_{3}^{-1}\end{bmatrix}\begin{bmatrix}JX^{T}A\\ (X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}\end{bmatrix}$
$=\dfrac{1}{\rho}AXX^{T}A+X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}\Gamma_{3}^{-1}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}.$ (13)
In the framework of the tractable metric whose representing matrix has just been obtained, the targeted generalized canonical metric applied to two tangent vectors $Z_{i}=XW_{i}+A^{-1}X_{\perp}K_{i}$, $i=1,2,$ at any $X\in\mathrm{iSt}(k,n)$ is
$g_{\rho,X_{\perp}}(Z_{1},Z_{2})\coloneqq\mathrm{tr}\left(Z_{1}^{T}\mathbf{M}_{X}Z_{2}\right)=\dfrac{1}{\rho}\mathrm{tr}(W_{1}^{T}W_{2})+\mathrm{tr}(K_{1}^{T}\Gamma_{3}^{-1}K_{2}).$ (14)
Obviously, $g_{\rho,X_{\perp}}$ is a non-degenerate bilinear form defined on $T_{X}\mathrm{iSt}(k,n)$ and dependent upon both $\rho$ and $X_{\perp}$. To make this a Riemannian metric, except for the special parameter $\rho$, it is essential to require that $\mathbf{M}_{X}$ depends only on $X\in\mathrm{iSt}(k,n)$ and this dependence is smooth. To this end, we will impose additional conditions on $X_{\perp}$ in terms of $\Gamma_{3}$. If we choose
$\left(X_{\perp}^{T}X_{\perp}\right)^{-1}=\Gamma_{3}$ (15)
then the second term of (13) becomes
$X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}\left(X_{\perp}^{T}X_{\perp}\right)\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}=\left(A(I_{n}-XJX^{T}A)\right)^{2}$
thanks to the identity (5). Thus, the representing matrix is
$\mathbf{M}_{X}=\frac{1}{\rho}AXX^{T}A+\left(A\left(I_{n}-XJX^{T}A\right)\right)^{2}.$ (16)
Alternatively, we can take
$\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}=\Gamma_{3}.$ (17)
Accordingly, it follows that
$X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}\Gamma_{3}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}$ $=X_{\perp}\left(X_{\perp}^{T}X_{\perp}\right)^{-1}X_{\perp}^{T}$
$=I_{n}-X\left(X^{T}X\right)^{-1}X^{T}.$
Thus,
$\mathbf{M}_{X}=\frac{1}{\rho}AXX^{T}A+I_{n}-X\left(X^{T}X\right)^{-1}X^{T}.$
Remark 1
  • One can see that when $\rho=2$ and $\Gamma_{3}=I_{n-k}$, the formulation (16) reduces to the canonical metric given in (EdelAS98, , Eq. (2.39)). Hence, the name generalized canonical metric. Moreover, we use this value for $\rho$ in numerical tests unless stated otherwise.
  • The fact that the generalized canonical metric (14) is independent of $X_{\perp}$ upon the choices for $\Gamma_{3}$ can also be derived using (4). Indeed, the dependence of the metric on $X_{\perp}$ is only on the second term which is now eliminated by using either (15) or (17). Namely, for the first choice,
    $K^{T}_{1}\Gamma_{3}^{-1}K_{2}=K^{T}_{1}X_{\perp}^{T}X_{\perp}K_{2}=\left(X_{\perp}K_{1}\right)^{T}X_{\perp}K_{2}$
    is independent of $X_{\perp}$ because $A^{-1}X_{\perp}K_{j}$ is so for $j=1,2,$ according to (4). For the second one, we obtain that
    $K^{T}_{1}\Gamma_{3}^{-1}K_{2}$ $=K^{T}_{1}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)\left(X_{\perp}^{T}X_{\perp}\right)^{-1}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)K_{2}$
    $=Z_{1}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)\left(X_{\perp}^{T}X_{\perp}\right)^{-1}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)$
    $\qquad\qquad\qquad\times\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}Z_{2}$
    $=Z_{1}^{T}X_{\perp}\left(X_{\perp}^{T}X_{\perp}\right)^{-1}X_{\perp}^{T}Z_{2}$
    $=Z_{1}^{T}\left(I_{n}-X\left(X^{T}X\right)^{-1}X^{T}\right)Z_{2}.$
3.1 Geometry under the generalized canonical metric
For any $X\in\mathrm{iSt}(k,n),$ the matrix $\mathbf{M}_{X}$ defined by (13) is symmetric positive-definite, so we can extend the inner product $g_{\rho,X_{\perp}}$ on $T_{X}\mathrm{iSt}(k,n)$ to an inner product on $\mathbb{R}^{n\times k}$} which is given by $g_{\rho,X_{\perp}}(Z_{1},Z_{2})=\mathrm{tr}(Z_{1}^{T}\mathbf{M}_{X}Z_{2}),$ for all $Z_{1},Z_{2}\in{\mathbb{R}^{n\times k}}.$ The normal space to $\mathrm{iSt}(k,n)$ at $X$ with respect to $g_{\rho,X_{\perp}}$ is defined by
$T_{X}\mathrm{iSt}(k,n)_{\rho}^{\perp}\coloneqq\left\{N\in{\mathbb{R}^{n\times k}}:g_{\rho,X_{\perp}}(N,Z)=0\text{ for all }Z\in T_{X}\mathrm{iSt}(k,n)\right\}.$
The following statement characterizes the elements of this space. Although the metric (13) depends on $X_{\perp}$, the elements of the normal space, as we will see, do not.
Proposition 1
$T_{X}\mathrm{iSt}(k,n)_{\rho}^{\perp}=\left\{XU:JU\in\mathrm{Sym}(k)\right\}.$
Proof Let $N=XU_{N}+A^{-1}X_{\perp}K_{N}$ with $U_{N}\in{\mathbb{R}^{k\times k}},$ $K_{N}\in\mathbb{R}^{(n-k)\times k}$ as in (4). For each $Z\in T_{X}\mathrm{iSt}(k,n),$ we have $Z=XW_{Z}+A^{-1}X_{\perp}K_{Z},$ with $JW_{Z}\in\mathrm{Skew}(k),K_{Z}\in\mathbb{R}^{(n-k)\times k}$ as in (6). By (14), we have
$g_{\rho,X_{\perp}}(N,Z)$ $=\dfrac{1}{\rho}\mathrm{tr}(U_{N}^{T}W_{Z})+\mathrm{tr}(K_{N}^{T}\Gamma_{3}^{-1}K_{Z})$
$=\dfrac{1}{\rho}\mathrm{tr}\left[(JU_{N})^{T}(JW_{Z})\right]+\mathrm{tr}(K_{N}^{T}\Gamma^{-1}_{3}K_{Z}).$
Therefore, $g_{\rho,X_{\perp}}(N,Z)=0$, for any $JW_{Z}\in\mathrm{Skew}(k)$ and any $K_{Z}\in\mathbb{R}^{(n-k)\times k},$ if and only if $K_{N}=0$ and $JU_{N}\in\mathrm{Sym}(k)$ since $\mathrm{Skew}(k)$ is the orthogonal complement of $\mathrm{Sym}(k)$ with respect to the Euclidean product on $\mathbb{R}^{n\times k}$}. Therefore, $N=XU_{N}$ with $JU_{N}\in\mathrm{Sym}(k)$ if and only if $N\in T_{X}\mathrm{iSt}(k,n)_{\rho}^{\perp}$ which is also the expected characterization. ∎
Next, we present formulae for the orthogonal projections $\mathscr{P}_{X}$ and $\mathscr{P}_{X}^{\perp}$ with respect to the inner product $g_{\rho,X_{\perp}}$ onto $T_{X}\mathrm{iSt}(k,n)$ and $T_{X}\mathrm{iSt}(k,n)^{\perp}_{\rho},$ respectively.
Proposition 2
Given any $Y\in{\mathbb{R}^{n\times k}},$ the orthogonal projections of $Y$ onto $T_{X}\mathrm{iSt}(k,n)$ and $T_{X}\mathrm{iSt}(k,n)^{\perp}_{\rho}$ are respectively given by
$\mathscr{P}_{X}(Y)$ $=Y-XJ\mathrm{sym}(X^{T}AY)$ (18)
$=XJ\mathrm{skew}(X^{T}AY)+(I_{n}-XJX^{T}A)Y$ (19)
and
$\mathscr{P}_{X}^{\perp}(Y)=XJ\mathrm{sym}(X^{T}AY).$ (20)
Proof From (6) and Proposition 1, the images of $Y$ should take the forms $\mathscr{P}_{X}(Y)=XW_{Y}+A^{-1}X_{\perp}K_{Y}$ and $\mathscr{P}_{X}^{\perp}(Y)=XU_{Y},$ in which $JW_{Y}\in\mathrm{Skew}(k),$ $K_{Y}\in\mathbb{R}^{(n-k)\times k},$ and $JU_{Y}\in\mathrm{Sym}(k).$ In view of (7), it holds that
$Y=XW_{Y}+A^{-1}X_{\perp}K_{Y}+XU_{Y}.$ (21)
Left multiplying this identity by $X^{T}A$ yields $X^{T}AY=JW_{Y}+JU_{Y}$. Therefore, $\mathrm{sym}(X^{T}AY)=JU_{Y}$ leads to $\mathscr{P}_{X}^{\perp}(Y)=XU_{Y}=XJ\mathrm{sym}(X^{T}AY)$ which proves (20).
The identity (18) follows from (21),
$\mathscr{P}_{X}(Y)=XW_{Y}+A^{-1}X_{\perp}K_{Y}=Y-XU_{Y}=Y-XJ\mathrm{sym}(X^{T}AY).$
Since $\mathrm{sym}(X^{T}AY)=\mathrm{skew}(X^{T}AY)-X^{T}AY,$ (19) is a consequence of (18). ∎
3.2 Riemannian gradient
We discuss now the Riemannian gradient of the cost function $f$ at $X\in\mathrm{iSt}(k,n)$ with respect to the generalized canonical metric $g_{\rho,X_{\perp}}$ constructed in Section 3. Recalling that it is the unique vector of $T_{X}\mathrm{iSt}(k,n)$, denoted by $\mathrm{grad}_{gc}f(X)$, satisfying
$g_{\rho,X_{\perp}}(\mathrm{grad}_{gc}f(X),Z)=\textsc{D}\bar{f}(X)[Z]=\langle\nabla\bar{f}(X),Z\rangle\text{ for all }Z\in T_{X}\mathrm{iSt}(k,n),$
where $\bar{f}$ is any smooth extension of $f$ around $X$ in $\mathbb{R}^{n\times k}$}. We have the following formula of $\mathrm{grad}_{gc}f(X).$
Proposition 3
The Riemannian gradient of the cost function $f$ at $X$ in $\mathrm{iSt}(k,n)$ is given by
$\mathrm{grad}_{gc}f(X)=\rho\,XJ\mathrm{skew}(JX^{T}\nabla\bar{f}(X))+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}\nabla\bar{f}(X).$ (22)
Proof In view of the definition of the Riemannian gradient and (14), there holds
$\langle\mathbf{M}_{X}\mathrm{grad}_{gc}\bar{f}(X),Z\rangle=g_{\rho,X_{\perp}}\left(\mathrm{grad}_{gc}\bar{f}(X),Z\right)=\textsc{D}\bar{f}(X)[Z]=\langle\nabla\bar{f}(X),Z\rangle,$
for any $Z\in{\mathbb{R}^{n\times k}}.$ Therefore, from (12),
$\mathrm{grad}_{gc}\bar{f}(X)=\mathbf{M}_{X}^{-1}\nabla\bar{f}(X)=\left(\rho XX^{T}+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}\right)\nabla\bar{f}(X).$
It follows from (AbsiMS08, , Eq. (3.37)) and (19) that
$\mathrm{grad}_{gc}f(X)$ $=\mathscr{P}_{X}(\mathrm{grad}_{gc}\bar{f}(X))$
$=XJ\mathrm{skew}\biggr(X^{T}A\left(\rho XX^{T}+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}\right)\nabla\bar{f}(X)\biggr)$
$\quad+\;(I_{n}-XJX^{T}A)\left(\rho XX^{T}+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}\right)\nabla\bar{f}(X)$
$=\rho XJ\mathrm{skew}\left(JX^{T}\nabla\bar{f}(X)\right)+\left(\rho XX^{T}+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}\right)\nabla\bar{f}(X)$
$\quad-\rho XX^{T}\nabla\bar{f}(X)$
$=\rho XJ\mathrm{skew}\left(JX^{T}\nabla\bar{f}(X)\right)+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}\nabla\bar{f}(X).$
This proves the identity (22). ∎
It is stressed that in practice, the appearance of $X_{\perp}$ in (22) needs to be removed. This must be made in accordance with the condition imposed in either (15) or (17). Consider first the case $\left(X_{\perp}^{T}X_{\perp}\right)^{-1}=\Gamma_{3}$. It is readily checked that
$X_{\perp}\Gamma_{3}X_{\perp}^{T}=X_{\perp}\left(X_{\perp}^{T}X_{\perp}\right)^{-1}X_{\perp}^{T}=I_{n}-X(X^{T}X)^{-1}X^{T}.$
Hence, the expression (22) becomes
$\mathrm{grad}_{gc}f(X)=$ $\rho\,XJ\mathrm{skew}(JX^{T}\nabla\bar{f}(X))$ (23)
$+A^{-1}\left(I_{n}-X(X^{T}X)^{-1}X^{T}\right)A^{-1}\nabla\bar{f}(X).$
In the case that $\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}=\Gamma_{3}$, taking (5) into account, it follows that
$A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}$ $=A^{-1}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}A^{-1}$
$=\left(I_{n}-XJX^{T}A\right)\left(I_{n}-XJX^{T}A\right)^{T}.$
Thus, we obtain
$\mathrm{grad}_{gc}f(X)=$ $\rho\,XJ\mathrm{skew}(JX^{T}\nabla\bar{f}(X))$ (24)
$+\left(I_{n}-XJX^{T}A\right)\left(I_{n}-XJX^{T}A\right)^{T}\nabla\bar{f}(X).$
If matrix $A$ is dense and has no special structure, the computational cost for $A^{-1}(\cdot)$ in (23) is generally comparable to $n^{3}$. Therefore in this case, using (24) is more advantageous. One can see that the computation of (24) requires $2n^{2}k+10nk^{2}-nk-3k^{2}$ flops. Obviously, when $k$ is approaching $n$, this cost is cheaper than that caused by (9). This will be seen in the numerical examples presented in Section 6.
Before closing this section, we discuss the reason for some closed-form expressions of the orthogonal projections in popular instances. Indeed, our general setting here and the tractable metric in (TiepSon2025, , Sect. 3) allow a clear explanation. First, for the orthogonal Stiefel manifold, the canonical metric proposed in (EdelAS98, , Subsect. 2.3) is a special case of our proposed metric (14) when $\rho=2$ and $\Gamma_{3}=I_{n-k}$. Therefore, there is no surprise when the associated orthogonal projection on the tangent space is in closed-form; namely, (EdelAS98, , Eq. (2.4)) is a special instance of (19). Alternatively, if the Euclidean metric is used, i.e., $\mathbf{M}_{X}=I_{n}$, equation (8) immediately admits the explicit solution $U_{X,Y}=\mathrm{sym}(X^{T}Y)$, which together with (TiepSon2025, , Prop. 3.2), yields (EdelAS98, , Eq. (2.4)). Second, for the generalized Stiefel manifold, it was suggested in ShusA2023 ; WangDPY2024 to choose $\mathbf{M}_{X}=A$. This results in the explicit solution $U_{X,Y}=\mathrm{sym}(X^{T}AY)$ to (8). In view of this, (TiepSon2025, , Prop. 3.2) implies (ShusA2023, , Eq. (3.18)) and (WangDPY2024, , Eq. (4.3)). Note, however, that unlike our derivation, (ShusA2023, , Eq. (3.18)) was obtained from (EdelAS98, , Eq. (2.4)) via a change of variable.
This explanation is also applicable to the geometry of the symplectic Stiefel manifold under the canonical-like metric GSAS21 as a special case of the tractable metric GSS2024_2 .
4 Quasi-geodesics with respect to the generalized canonical metric
A geodesic on a manifold is a smooth curve with vanishing acceleration. In the dawn of Riemannian optimization, the solution is updated by searching along the geodesic which goes through the current solution in an appropriate tangent direction EdelAS98 ; Gaba82 ; Smit94 . In the Euclidean space, zero acceleration means $\ddot{X}(t)=0$ which implies that $X(t)=\dot{X}(0)t+X(0)$ which is a straight line. That is, geodesics can be viewed as a generalization of lines in the Euclidean space to the Riemannian manifold. In a general setting, the notion of acceleration depends on the Riemannian connection (AbsiMS08, , Sect. 5.4) and therefore the differential equation determining geodesics is usually much more involved. This makes the task of finding an explicit formula for the geodesic, given initial information $X(0),\dot{X}(0)$, very difficult. Fortunately, in optimization, only local information encoded by geodesics and its feasibility are interesting due to their usage in updating the iterate. Thus, one usually retains these features in approximating geodesics which results in the so-called quasi-geodesics, see, e.g., JurdML2020 ; KrakMLB2017 ; NishA2005 for more approaches to the quasi-geodesics on the Stiefel manifold.
In this paper, we follow a procedure similar to that of (GSAS21, , Subsect. 4.3), which was adopted from (EdelAS98, , Subsect. 2.2). Namely, we replace the acceleration that should be formulated in accordance with the Riemannian connection associated with the proposed generalized canonical metric by the one corresponding to the Euclidean metric, which turns out to be $\ddot{Y}(t).$ In view of the normal vectors formulated in Proposition 1, we define the quasi-geodesic with respect to the generalized canonical metric $g_{\rho,X_{\perp}}$ to be any solution $Y_{\rho}(t)\in\mathrm{iSt}(k,n)$ of the initial value problem
$\ddot{Y}(t)+Y(t)U(t)=0,\ Y(0)=X\in\mathrm{iSt}(k,n),\ \dot{Y}(0)=Z\in T_{X}\mathrm{iSt}(k,n),$ (25)
with some $U(t)$ such that $JU\in\mathrm{Sym}(k)$ for all $t\geq 0$, although the obtained result apparently holds for $t<0$ too. In this section, a closed-form solution to this equation is presented. For the sake of brevity, we refrain from writing the variable $t$ in our presentation.
As the first step, we show that for the equation (25) to have a solution, $U(t)$ must have a special form. Consecutively differentiating with respect to $t$ both sides of the identity $Y^{T}AY=J$ twice yield
$\dot{Y}^{T}AY+Y^{T}A\dot{Y}$ $=0,$ (26)
$\ddot{Y}^{T}AY+2\dot{Y}^{T}A\dot{Y}+Y^{T}A\ddot{Y}$ $=0,$ (27)
respectively. Substituting $\ddot{Y}=-YU$ from (25) into (27) gives
$(-YU)^{T}AY+2\dot{Y}^{T}A\dot{Y}+Y^{T}A(-YU)=0$ (28)
or $\dot{Y}^{T}A\dot{Y}=JU$ due to the symmetry of $JU$. Hence, $U(t)=J\dot{Y}^{T}A\dot{Y}$ and therefore, the equation (25) becomes
$\ddot{Y}+YJ(\dot{Y}^{T}A\dot{Y})=0.$ (29)
Next, let us set $W\coloneqq Y^{T}A\dot{Y}$ and $V\coloneqq\dot{Y}^{T}A\dot{Y}$. Thanks to the relation (26), $W\in\mathrm{Skew}(k)$ for $t\geq 0.$ Moreover, using (29), we get
$\dot{W}=V+Y^{T}A\ddot{Y}=V-Y^{T}AYJV=V-V=0,\mbox{ for }t\geq 0.$
Hence, $W(t)=W(0)\eqqcolon W_{0}\in\mathrm{Skew}(k)$ for $t\geq 0$. We proceed to determine $V$ which is symmetric. Taking the derivative of $V$ with (29) in mind, we have that
$\dot{V}$ $=\ddot{Y}^{T}A\dot{Y}+\dot{Y}^{T}A\ddot{Y}=(-YJV)^{T}A\dot{Y}+\dot{Y}^{T}A(-YJV)$
$=-VJW+WJV=W_{0}JV-VJW_{0}.$
We will show that the initial value problem
$\dot{V}=W_{0}JV-VJW_{0},\ V(0)=V_{0},$ (30)
admits a solution $V(t)=\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}.$ Indeed, differentiating such $V(t)$ yields
$\dot{V}$ $=-\mathrm{e}^{-tW_{0}^{T}J}W_{0}^{T}JV_{0}\mathrm{e}^{-tJW_{0}}-\mathrm{e}^{-tW_{0}^{T}J}V_{0}JW_{0}\mathrm{e}^{-tJW_{0}}$
$=-W_{0}^{T}J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}-\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}JW_{0}$
$=-W_{0}^{T}JV-VJW_{0}=W_{0}JV-VJW_{0}.$
Here, we invoke the skew-symmetry of $W_{0}$ and the identity $M\mathrm{e}^{tM}=\mathrm{e}^{tM}M$ for any real square matrix $M$ and any real number $t$.
Next, we present a technical lemma, which is similar to (GSAS21, , Lem. 4.6).
Lemma 1
If $\Theta\in\mathrm{Skew}(k)$ and $\Gamma\in\mathrm{Sym}(k),$ then for any $t\in\mathbb{R},$ it holds that
$\left(\mathrm{exp}\left(t\begin{bmatrix}J\Theta&J\Gamma\\ I_{k}&J\Theta\end{bmatrix}\right)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\mathrm{exp}\left(t\begin{bmatrix}J\Theta&J\Gamma\\ I_{k}&J\Theta\end{bmatrix}\right)=\begin{bmatrix}0&J\\ -J&0\end{bmatrix}$. (31)
Proof For the sake of convenience, let us set $\Lambda\coloneq\left[\begin{smallmatrix}J\Theta&J\Gamma\\ I_{k}&J\Theta\end{smallmatrix}\right]$ and the function on the left hand side of (31) as $u(t)\coloneq\left(\mathrm{exp}\left(t\Lambda\right)\right)^{T}\left[\begin{smallmatrix}0&J\\ -J&0\end{smallmatrix}\right]\mathrm{exp}\left(t\Lambda\right)$. Then, it follows that
$\dot{u}(t)=\left(\mathrm{exp}\left(t\Lambda\right)\right)^{T}\biggr(\Lambda^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}+\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Lambda\biggr)\mathrm{exp}\left(t\Lambda\right).$
By assumption, one can easily check that
$\Lambda^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}+\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Lambda=0.$
Therefore, $\dot{u}(t)=0$ which leads to the fact that $u(t)=u(0)=\begin{bmatrix}0&J\\ -J&0\end{bmatrix}$, for any $t\in\mathbb{R}.$
Now, we are ready to give the explicit formula for a solution to the problem (25).
Proposition 4
Given $X\in\mathrm{iSt}(k,n),Z\in T_{X}\mathrm{iSt}(k,n).$ The initial value problem
$\ddot{Y}(t)+Y(t)J\dot{Y}(t)^{T}A\dot{Y}(t)=0,\ Y(0)=X,\ \dot{Y}(0)=Z,$ (32)
has a unique global solution in $\mathrm{iSt}(k,n)$ given by
$Y_{\rho}(t)=Y_{\rho}^{\mathrm{qgeo}}(t,Z)\coloneqq[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$ (33)
where $W_{0}=Y_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=X^{T}AZ$ and $V_{0}\coloneq\dot{Y}_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=Z^{T}AZ.$
Proof First, let us assume that the problem (32) has a global solution in $\mathrm{iSt}(k,n),$ denoted also by $Y_{\rho}(t)$. It follows that
$\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\right)$ $=\dot{Y_{\rho}}(t)\mathrm{e}^{tJW_{0}}+Y_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}.$
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right)$ $=\ddot{Y_{\rho}}(t)\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
$=\left(-Y_{\rho}(t)JV\right)\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
$=-Y_{\rho}(t)J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
$=-Y_{\rho}(t)\mathrm{e}^{tJW_{0}}JV_{0}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0},$
in which we have invoked (30) to replace $V\coloneq\dot{Y_{\rho}}^{T}A\dot{Y_{\rho}}$ by its explicit form $V=\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}$ when $Y_{\rho}(t)\in\mathrm{iSt}(k,n),$ for all $t\geq 0.$ Therefore,
$\frac{\mathrm{d}}{\mathrm{d}t}\left[Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right]=\left[Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right]\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}$. (34)
This basic equation admits an explicit solution from which we obtain that
$Y_{\rho}(t)$ $=\left[Y_{\rho}(0)\;\;\dot{Y}_{\rho}(0)\right]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$
which is the expected formula.
Next, we will show that this is indeed a global solution to the problem (32). For brevity, we set $\Psi\coloneq\left[\begin{smallmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{smallmatrix}\right]$. In view of (33), it holds that
$\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}+Y_{\rho}\mathrm{e}^{tJW_{0}}JW_{0}=\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}\mathrm{e}^{tJW_{0}}\right)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}$.
This leads to
$\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}-[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JW_{0}$ (35)
$=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\biggr(\begin{bmatrix}JW_{0}\\ I_{k}\end{bmatrix}-\begin{bmatrix}JW_{0}\\ 0\end{bmatrix}\biggr)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
Repeating this procedure with $\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}$, we obtain
$\ddot{Y}_{\rho}\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}JW_{0}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
Using (35), this identity gives
$\ddot{Y}_{\rho}\mathrm{e}^{tJW_{0}}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}-JV_{0}\\ 0\end{bmatrix}$. (36)
On the other hand, using the expression for $\dot{Y}_{\rho}$ in (35), we receive the identities
$\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho}=$ $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}},$ (37)
in which Lemma 1 have been invoked in the fifth identity.
Finally, left multiplying (37) by $Y_{\rho}J$ yields
$Y_{\rho}J(\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho})=$ $\,Y_{\rho}J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}J^{2}\mathrm{e}^{tJW_{0}}JV_{0}\mathrm{e}^{-tJW_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JV_{0}\mathrm{e}^{-tJW_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}JV_{0}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$ (38)
in which the second identity is due to the fact that
$(\mathrm{e}^{-tJW_{0}})^{T}=\mathrm{e}^{-tW_{0}^{T}J}=J^{2}\mathrm{e}^{-tW_{0}^{T}J}=J\mathrm{e}^{tJW_{0}}J.$
The conclusion follows from the two identities (36) and (38).
Finally, the uniqueness of the solution is derived from that of the problems (30) and (34) which is a well-known fact in the theory of linear differential equations in Banach spaces, see, e.g., (Deimling1977, , Sect. 1.1). ∎
The second part of the proof for Proposition 4, implicitly assumes that the curve $Y_{\rho}(t)$ in (33) is feasible. The statement below confirms it.
Proposition 5
The curve $Y_{\rho}(t)$ given in (33) lies in the indefinite Stiefel manifold $\mathrm{iSt}_{A,J}(k,n)$, i.e., for any $t\geq 0,$ we have $Y_{\rho}(t)^{T}AY_{\rho}(t)=J.$
Proof From (33), it follows that
$Y_{\rho}^{T}AY_{\rho}=\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[I_{k}\;\;0]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}.$ (39)
According to the proof for (37), we have that
$\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)=\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi=\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}$.
This fact leads (39) to
$Y_{\rho}^{T}AY_{\rho}=$ $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[I_{k}\;\;0]\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $J\mathrm{e}^{tJW_{0}}J^{2}\mathrm{e}^{-tJW_{0}}=J,$
in which the second identity follows from the fact that $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}=J\mathrm{e}^{tJW_{0}}J.$
In the case of generalized Stiefel manifold, i.e., $A$ is spd and $J=I_{k}$, the obtained quasi-geodesic reduces to a simple form as follows.
Corollary 1
Given $X\in\mathrm{St}_{A}(k,n)$, $Z\in T_{X}\mathrm{St}_{A}(k,n)$ and set $W_{0}=X^{T}AZ$, $V_{0}=Z^{T}AZ$. The curve
$Y(t)=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}W_{0}&-V_{0}\\ I_{k}&W_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tW_{0}}$ (40)
is a quasi-geodesic associated with $X$ and $Z$ in the sense that $Y(t)\in\mathrm{St}_{A}(k,n)$ for all $t\geq 0$ and $Y(0)=X,\dot{Y}(0)=Z$.
Remark 2
Restricting to the orthogonal Stiefel manifold, the quasi-geodesic (40) reduces to the geodesic obtained in (EdelAS98, , Subsect. 2.2). It is worth to note that the one in (EdelAS98, , Subsect. 2.2) is indeed associated with the Euclidean metric. Nevertheless, by direct verification, one can show that the normal space, orthogonal projections onto the tangent and normal spaces under the Euclidean and under the canonical metrics are identical. Therefore, for the orthogonal Stiefel manifold, the embedded geodesic associated with the Euclidean metric is also the quasi-geodesic associated with the canonical metric. Hence, (33) is a generalization of this quasi-geodesic to the indefinite Stiefel manifold.
5 Quasi-geodesic based retraction on $\mathrm{iSt}(k,n)$
Let us first formally define a retraction on $\mathrm{iSt}(k,n)$ AbsiMS08 ; AdleDMMS02 . It is a smooth mapping $\mathcal{R}$ from the tangent bundle $T\mathrm{iSt}(k,n)\coloneqq\bigcup_{X\in\mathrm{iSt}(k,n)}T_{X}\mathrm{iSt}(k,n)$ to $\mathrm{iSt}(k,n)$ satisfying the following conditions for all $X\in\mathrm{iSt}(k,n):$
$\mathcal{R}_{X}(0)=X,\text{ where }0\text{ is the origin of }T_{X}\mathrm{iSt}(k,n),$ (R.1)
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{R}_{X}\left(tZ\right)\rvert_{t=0}=Z,\text{ for all }Z\in T_{X}\mathrm{iSt}(k,n),$ (R.2)
where $\mathcal{R}_{X}$ is the restriction of $\mathcal{R}$ to $T_{X}\mathrm{iSt}(k,n).$ The crucial role of a retraction in optimization is to maintain the feasibility of the updated iterate from a feasible search direction which is a tangent vector at the current iterate. In TiepSon2025 , a retraction based on the Cayley transformation was proposed. In this work, based on the formulation of a quasi-geodesic (33), we propose another retraction which will be termed quasi-geodesic retraction as follows:
$\mathcal{R}^{\mathrm{qgeo}}:T\mathrm{iSt}(k,n)$ $\longrightarrow\mathrm{iSt}(k,n)$
$Z\in T_{X}\mathrm{iSt}(k,n)$ $\longmapsto\mathcal{R}_{X}^{\mathrm{qgeo}}(Z)\coloneqq Y_{\rho}^{\mathrm{qgeo}}(1,Z)$ (41)
$=[X\;\;Z]\mathrm{exp}\left(\begin{bmatrix}JX^{T}AZ&-JZ^{T}AZ\\ I_{k}&JX^{T}AZ\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-JX^{T}AZ},$
where $\mathcal{R}_{X}^{\mathrm{qgeo}}$ is the restriction of $\mathcal{R}^{\mathrm{qgeo}}$ to $T_{X}\mathrm{iSt}(k,n).$ The well-definedness of $\mathcal{R}^{\mathrm{qgeo}}$ is guaranteed in the following result whose proof is similar to (GSAS21, , Lem. 5.1).
Proposition 6
The map $\mathcal{R}^{\mathrm{qgeo}}$ given in (41) is a globally defined retraction on $\mathrm{iSt}(k,n).$
Proof We only need to verify the conditions (R.1) and (R.2) since the expression in (41) is well-defined for any $X\in\mathrm{iSt}(k,n)$ and $Z\in T_{X}\mathrm{iSt}(k,n)$ thanks to Proposition 5.
It holds that $\mathcal{R}_{X}^{\mathrm{qgeo}}(0)=[X\;\;0]\mathrm{exp}\left(\left[\begin{smallmatrix}0&0\\ I_{k}&0\end{smallmatrix}\right]\right)\left[\begin{smallmatrix}I_{k}\\ 0\end{smallmatrix}\right]\mathrm{e}^{0}=[X\;\;0]\left[\begin{smallmatrix}I_{k}\\ I_{k}\end{smallmatrix}\right]=X$. Moreover, for any number $t\geq 0,$
$\mathcal{R}_{X}^{\mathrm{qgeo}}(tZ)=[X\;\;tZ]\mathrm{exp}\left(\begin{bmatrix}tJX^{T}AZ&-t^{2}JZ^{T}AZ\\ I_{k}&tJX^{T}AZ\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJX^{T}AZ}.$
So,
[…]
$+[X\;\;0]\mathrm{exp}\left(\begin{bmatrix}0&0\\ I_{k}&0\end{bmatrix}\right)\begin{bmatrix}JX^{T}AZ&0\\ 0&JX^{T}AZ\end{bmatrix}\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{0}$
$-[X\;\;0]\mathrm{exp}\left(\begin{bmatrix}0&0\\ I_{k}&0\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{0}JX^{T}AZ$
$=Z,$
where we have invoked the identity $\frac{\mathrm{d}t}{t}\mathrm{e}^{M(t)}|_{t=0}=\mathrm{e}^{M(0)}\dot{M}(0)$ for smooth one-parameter matrix $M(t)$ such that $M(0)\dot{M}(0)=\dot{M}(0)M(0)$, see (Wilcox1967, , Eq. (2.1)). ∎
6 Numerical examples
In this section, theoretical results are assembled to enable a Riemannian gradient descent scheme and then numerically verified via examples.
A Riemannian gradient descent method
To minimize a continuously differentiable cost function $f$ on the indefinite Stiefel manifold $\mathrm{iSt}_{A,J}$,
$\min_{X\in\mathbb{R}^{n\times k}}f(X)\ \mbox{ s.t. }\ X^{T}AX=J,$ (42)
given an initial point $X_{0}\in\mathrm{iSt}_{A,J}(k,n)$, one consecutively updates the approximate solution by
$X_{j+1}=\mathcal{R}_{X_{j}}(-\tau_{j}\mathrm{grad}f(X_{j})),$
where $\mathcal{R}_{X_{j}}(\cdot)$ is a retraction at $X_{j}$, $\mathrm{grad}f(X_{j})$ denotes the Riemannian gradient of the cost function $f$ at $X_{j}$ with respect to a chosen Riemannian metric $\mathrm{g}(\cdot,\cdot)$, and $\tau_{j}>0$ is a step size to be determined. To compute the step size, we combine a nonmonotone line search technique with the alternating Barzilai-Borwein strategy on an embedded submanifold, see BarB88 ; HuLWY2020 ; IannP18 ; ZhanH2004 . For the sake of convenience, we recall the details in Algorithm 1. Note from (43) that if $\alpha$ is chosen to be zero, this scheme reduces to the popular monotonically decreasing line search with Amijo’s rule. Algorithm 1 was also employed in TiepSon2025 for numerical verification of optimization using the geometric quantities under a general tractable metric and the Cayley retraction. Its global convergence is guaranteed by (GSAS21, , Thm. 5.7). Namely, regardless of the initial point $X_{0}$, each accumulation point $X^{*}$ of the iterate generated by Algorithm 1 satisfies the first-order necessary optimality condition $\mathrm{grad}f(X^{*})=0$.
Algorithm 1 Riemannian gradient algorithm for the optimization problem (42)
0: The cost function $f$, initial guess $X_{0}\in\mathrm{iSt}_{A,J}$, $\gamma_{0}>0$, $0<\gamma_{\min}<\gamma_{\max}$, $\beta,$ $\delta\in(0,1)$, $\alpha\in[0,1]$, $q_{0}=1$, $c_{0}=f(X_{0})$.
0: Sequence of iterates $\{X_{j}\}$.
1:for $j=0,1,2,\dots$ do
2:  Compute $Z_{j}=-\mathrm{grad}f(X_{j})$.
3:  if $j>0$ then
4:    $\gamma_{j}=\left\{\begin{array}[]{ll}\dfrac{\langle W_{j-1},W_{j-1}\rangle}{\left|\mathrm{tr}(W_{j-1}^{T}Y_{j-1})\right|}&\text{for odd }j,\\[14.22636pt] \dfrac{\left|\mathrm{tr}(W_{j-1}^{T}Y_{j-1})\right|}{\langle Y_{j-1},Y_{j-1}\rangle}&\text{for even }j,\end{array}\right$.where $W_{j-1}=X_{j}-X_{j-1}$ and $Y_{j-1}=Z_{j}-Z_{j-1}$.
5:  end if
6:  Calculate the trial step size $\gamma_{j}=\max\bigl(\gamma_{\min},\min(\gamma_{j},\gamma_{\max})\bigr)$.
7:  Find the smallest integer $\ell$ such that the nonmonotone condition
$f\big(\mathcal{R}(\tau_{j}Z_{j})\big)\leq c_{j}+\beta\,\tau_{j}\,\mathrm{g}\big(\mathrm{grad}f(X_{j}),Z_{j}\big)$ (43)
holds, where $\tau_{j}=\gamma_{j}\,\delta^{\ell}$.
8:  Set $X_{j+1}=\mathcal{R}(\tau_{j}Z_{j})$.
9:  Update $q_{j+1}=\alpha q_{j}+1$ and ${c_{j+1}=\frac{\alpha q_{j}}{q_{j+1}}c_{j}+\frac{1}{q_{j+1}}f(X_{j+1})}$.
10:end for
Algorithmic parameter setting
For the backtracking search, we set $\beta=1\mathrm{e}{-4}$, $\delta=5\mathrm{e}{-1}$, $\gamma_{0}=1\mathrm{e}{-3}$, $\gamma_{\min}=1\mathrm{e}{-15}$, $\gamma_{\max}=1\mathrm{e}{+5}$. Moreover, $\alpha=0.85$ is used in the nonmonotone condition. A running is considered convergent when the norm of the Riemannian gradient of the cost function at the current iterate is relatively small compared to that at the initial guess $X_{0}$, namely, $\|\mathrm{grad}f(X_{j})\|_{X_{j}}\leq\texttt{rstop}\|\mathrm{grad}f(X_{0})\|_{X_{0}}$. In all tests, rstop is set to be $1\mathrm{e}-5$.
Model problems
We examine two problems similar to that in TiepSon2025 . For the first one, we consider minimizing a trace cost function
$f(X):=\mathrm{tr}(X^{T}MX),\ X\in\mathrm{iSt}_{A,J}(k,n),$ (44)
where $n=1000,k=200$, $M=VV^{T}$ with $V$ is the $n\times(n-5)$ matrix randomly generated with default setting and then orthogonalized by orth in MATLAB. For the manifold, let $p=750,m=250,k_{p}=k_{m}=100$, we set $A=\mathrm{diag}(1,\ldots,p,-1,\ldots,-m)$ and $J=\mathrm{diag}(I_{k_{p}},-I_{k_{m}})$. The initial guess $X_{0}$ is set as the $n\times k$ zero matrix except for the submatrices $X_{0}(1,\ldots,k_{p};1,\ldots,k_{p})=\mathrm{diag}(\frac{1}{\sqrt{1}},\ldots,\frac{1}{\sqrt{k_{p}}})$ and $X_{0}(p+1,\ldots,p+k_{m};k_{p}+1,\ldots,k_{p}+k_{m})=\mathrm{diag}(\frac{1}{\sqrt{1}},\ldots,\frac{1}{\sqrt{k_{m}}})$.
The second model is the well-known Procrustes problem
$f(X)=\|GX-B\|^{2}_{F}$ (45)
on the $J$-orthogonal group with $n=500$, $J=\mathrm{diag}(I_{p},-I_{m}),$ where $p=375,$ $m=125$. Numerical data are generated as follow: $G$ is a $(n-20)\times n$ random matrix with default setting and $B$ is set as $G\mathrm{diag}(U,V)$ where $U\in\mathbb{R}^{p\times p}$ and $V\in\mathbb{R}^{m\times m}$ are two orthogonalized random matrices generated with the setting rng (1,’twister’). The $n\times n$ identity matrix is used as the initial guess.
Tests
It is worth to note that, while the Riemannian gradient is defined upon the metric, the retraction is metric-independent. Therefore, in Algorithm 1, one can combine any metric with any available retraction. Namely, there are the Cayley retraction and the one based on the quasi-geodesic developed in this paper. Regarding metrics, we have the general tractable and the generalized canonical metrics. In TiepSon2025 , the Euclidean and weighted Euclidean metrices have been used. The examples there showed that using the latter is much more advantageous. However, the Euclidean Hessian of the cost functions considered here are both degenerate; the choice of weighted Euclidean metric as in TiepSon2025 is inappropriate. Thus, we will use in our tests the Euclidean and the generalized canonical metrics. Consequently, there are totally four combinations whose corresponding numerical results will be marked by “Eucl-Cayley”, “Eucl-qgeo”, “gcan-Cayley”, and “gcan-qgeo”. Besides, since in our code we do not attempt to exploit the special structure of $A$, we always employ (17) for $X_{\perp}$. All the tests are done on a laptop with Intel(R) Core(TM) i7-4500U (at 1.8GHz, 5MB Cache, 8GB RAM) running MATLAB R2023a under Windows 10 Home. (The main code is made available at https://sites.google.com/view/ntson/code.)
We report the numerical results in Table 1 and Table 2 in which obj., grad., feas., #\#iter., #\#eval., and CPU represent the value of the cost function, the norm in the chosen metric of the Riemannian gradient of the cost function, the feasibility $\|X_{*}^{T}AX_{*}-J\|_{F}$, the number of iterations, the number of function evaluations, and the elapsed time, respectively. In the case that the iteration has not reached the stopping criterion, $X^{*}$ is replaced by the last iterate. In addition, the convergence histories are displayed in Figure 1 and Figure 2.
Table 1: Results for minimization of trace cost function (44).
Combination obj. grad. feas. #\#iter. #\#eval. CPU
Eucl-Cayley $0.6387$ $4.0204\mathrm{e}-6$ $2.3354\mathrm{e}-13$ 389 423 97.3941
Eucl-qgeo $0.6387$ $3.9532\mathrm{e}-6$ $5.0332\mathrm{e}-08$ 392 437 50.3922
gcan-Cayley 0.6387 $4.2671\mathrm{e}-6$ $2.5933\mathrm{e}-13$ 472 516 93.2189
gcan-qgeo 0.6387 $4.1930\mathrm{e}-6$ $3.1692\mathrm{e}-13$ 458 598 47.1033
Table 2: Results for minimization of Procrustes cost function (45).
Combination obj. grad. feas. #\#iter. #\#eval. CPU
Eucl-Cayley $21.5943$ $5.9039$ $2.5425\mathrm{e}-12$ 2000 2042 $2068$
Eucl-qgeo $38.1905$ $5.6773$ $8.4230\mathrm{e}-10$ 2000 2064 $2817$
gcan-Cayley 2.8671 $1.4755$ $3.3286\mathrm{e}-13$ 340 615 54
gcan-qgeo 15.5090 $1.5283$ $1.2650\mathrm{e}-10$ 402 738 356
Figure 1: Convergence history of the Riemannian gradient descent method with different combinations of metrics and retractions applied to the trace cost function (44).
Figure 2: Convergence history of the Riemannian gradient descent method with different combinations of metrics and retractions applied to the Procrustes cost function (45).
Discussion
Through examples presented, we can first see that, all newly established theoretical materials work properly and compatibly with the existing ones. Second, as expected, for the case that $k$ is not small, the schemes using the generalized canonical metric always need less time to be convergent (with respect to the given rstop) compared to the ones using the Euclidean metric; even in the case of Procrustes problem, within at most 2000 iterations, only the schemes with the generalized canonical metric converge. Third, the quasi-geodesic retraction apparently is less feasible than the Cayley retraction. This fact might come from the instability of the scaling and squaring algorithm invoked by expm in MATLAB which occasionally occurs as pointed out in GuetN2016 .
7 Conclusions
To avoid having to solve the Lyapunov matrix equation during the computation of Riemannian gradient for optimization on the indefinite Stiefel manifold, we proposed the generalized canonical metric and constructed the necessary associated material for running a Riemannian gradient descent scheme, including a retraction based on a quasi-geodesic corresponding to the new metric. The presented numerical examples verified the theoretical findings and suggested that, using the generalized canonical metric with the Cayley retraction is perhaps the most potential combination since it needs less time to reach a desired approximate solution and delivers better feasibility.
Acknowledgements.
This research was funded by the TNU-University of Sciences under the research group code: NNC.DHKH.2025.02.
Data availability
The authors confirm that the data supporting the findings of this study are available within the article.
Competing Interests
The authors declare that they have no conflict of interest.

### v2

@@
 Son, A Riemannian gradient descent method for optimization on the indefinite Stiefel manifold, arXiv:2410.22068v2[math.OC]], a tractable metric, which is indeed a family of Riemannian metrics defined by a symmetric positive-definite matrix depending on the contact point, has been used.
-In general, it requires solving a Lyapunov matrix equation every time when the gradient of the cost function is needed, which might significantly contribute to the computational cost.
+In general, it requires solving a Lyapunov matrix equation every time the gradient of the cost function is needed, which might significantly contribute to the computational cost.
 To address this issue, we propose a new Riemannian metric for the indefinite Stiefel manifold.
@@
 1 Introduction
-Optimization of smooth functions of matrix variable $X$ subject to the constraint of the form $\mathcal{M}=\{X\ :\ X^{T}AX=J\}$, where $A$ is nonsingular and symmetric and $J$ is symmetric, involutory, i.e., satisfying $J^{2}=I$, is an established line of researches due to its wide range of applications.
-Various problems such as dimensionality reduction of data via principal component analysis, canonical correlation analysis, finding locality preserving projection, and computation of the generalized eigenvalues of a positive-definite matrix pencil can be formulated as an optimization problem with such a constraint, see, e.g., HardSST04 ; HeN03 ; Hote92 ; KovaV95 ; SameW82 ; TiepSon2025 ; YgerBGR12 .
+Optimization of smooth functions of matrix variable $X$ subject to the constraint of the form $\mathcal{M}=\{X\ :\ X^{T}AX=J\}$, where $A$ is nonsingular and symmetric, and $J$ is symmetric, involutory, i.e., satisfying $J^{2}=I$, is an established line of research due to its wide range of applications.
+Various problems such as dimensionality reduction of data via principal component analysis, canonical correlation analysis, finding locality preserving projection, and computation of the generalized eigenvalues of a positive-definite matrix pencil can be formulated as an optimization problem with such a constraint; see, e.g., HardSST04 ; HeN03 ; Hote92 ; KovaV95 ; SameW82 ; TiepSon2025 ; YgerBGR12 .
 Additional conditions can be imposed on the matrices $A$ and $J$ upon the applications considered making the feasible set a differentiable manifold.
-When $A=I_{n}$ and $J=I_{k},k\leq n$, the $n\times n$ and $k\times k$ identity matrices, respectively, $\mathcal{M}$ is termed the (orthogonal) Stiefel manifold EdelAS98
+When $A=I_{n}$ and $J=I_{k},k\leq n$, the $n\times n$ and $k\times k$ identity matrices, respectively, $\mathcal{M}$ is termed the orthogonal Stiefel manifold EdelAS98
 $\mathrm{St}(k,n):=\{X\in{\mathbb{R}^{n\times k}}\ :\ X^{T}X=I_{k}\}.$
@@
 see, e.g., ShusA2023 .
-Although $\mathrm{St}(k,n)$ is a special case of $\mathrm{St}_{A}(k,n)$, one can see that $\mathrm{St}_{A}(k,n)$ can be turned into a Stiefel manifold by the mapping $X\mapsto R^{-1}Y$ where $A=R^{T}R$ is the Cholesky factorization of $A$.
+Although $\mathrm{St}(k,n)$ is a special case of $\mathrm{St}_{A}(k,n)$, one can see that $\mathrm{St}_{A}(k,n)$ can be turned into an orthogonal Stiefel manifold by the mapping $X\mapsto R^{-1}Y$ where $A=R^{T}R$ is the Cholesky factorization of $A$.
 A more general one in which $A$ is nonsingular, symmetric, and $J$ is symmetric, involutory, which covers the orthogonal and generalized Stiefel manifolds as special cases, is the indefinite Stiefel manifold
@@
 This manifold was recently introduced in TiepSon2025 .
-Exploiting the rich geometry of the feasible set, a popular approach, known as Riemannian optimization, is to extend well-known methods for unconstrained optimization in the Euclidean space to the case of these Riemannian manifolds once necessary geometric tools such as a Riemannian metric, orthogonal projections, a retraction are constructed.
+Exploiting the rich geometry of the feasible set, a popular approach, known as Riemannian optimization, is to extend well-known methods for unconstrained optimization in the Euclidean space to the case of these Riemannian manifolds once necessary geometric tools such as a Riemannian metric, an orthogonal projection, and a retraction are constructed.
 One is referred to classical works AdleDMMS02 ; EdelAS98 , monographs AbsiMS08 ; Boumal23 ; Sato2021 ; Udri1994 , and articles manopt ; MishS2016 ; SatoA19 ; ShusA2023 ; TiepSon2025 ; WangDPY2024 ; YgerBGR12 , to name a few among many others dedicated to optimization methods on the orthogonal, generalized, and indefinite Stiefel manifolds.
@@
 It not only helps quantify objects such as the length of vectors, the measure of angles but also has a great effect on the behavior of numerical methods.
-Specifically, it is a means of defining the Riemannian gradient of a cost function, a crucial material in most of optimization methods.
+Specifically, it is a means of defining the Riemannian gradient of a cost function, a crucial component in most of optimization methods.
 There are different ways to equip a manifold with a Riemannian metric depending on how the manifold is viewed.
-For some left/right-invariant Riemannian metrics, pseudo-Riemannian metrics equipped for the Stiefel manifold as a quotient manifold of two orthogonal groups, one is referred to (EdelAS98, , Subsect.
+For some left- or right-invariant Riemannian metrics, and pseudo-Riemannian metrics equipped for the orthogonal Stiefel manifold as a quotient manifold of two orthogonal groups, one is referred to (EdelAS98, , Subsect.
 2.3), JurdML2020 ; NishA2005 ; cf.
 BendZ21 for the symplectic Stiefel manifold.
-Here, we follow (TiepSon2025, , Prop.
-2.2) in which $\mathrm{iSt}_{A,J}(k,n)$ is shown to be an embedded submanifold of $\mathbb{R}^{n\times k}$} of dimension $nk-\frac{1}{2}k(k+1)$.
+Here, we follow TiepSon2025 in which $\mathrm{iSt}_{A,J}(k,n)$ is shown to be an embedded submanifold of $\mathbb{R}^{n\times k}$} of dimension $nk-\frac{1}{2}k(k+1)$.
 Then, a Riemannian metric is defined via an $n\times n$ spd $X$-smoothly depending matrix $\mathbf{M}_{X}$ in which the inner product of $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}_{A,J}(k,n)$ is determined as
@@
 3.3), which can be computationally expensive when $k$ is not small.
-It is surprising that in the case of Stiefel manifold, the Riemannian gradient with respect to the Euclidean and canonical metrics have closed-form expressions, see (EdelAS98, , Sect.
+It is surprising that in the case of the orthogonal Stiefel manifold, the Riemannian gradient with respect to the Euclidean and canonical metrics has closed-form expressions; see (EdelAS98, , Sect.
 2) and (AbsiMS08, , Subsect.
@@
 Note, however, that the motivation for using the canonical(-like) metric in EdelAS98 ; GSAS21 is to balance the contribution of the two components of tangent vectors in the metric formulation instead of obviating the necessity of solving the Lyapunov matrix equation as that for the metric proposed in this work.
-On the other hand, they inspire several questions: (1) How are we on track for not having to solve the Lyapunov matrix equations to compute the Riemannian gradient, (2) How to explain the different situations for orthogonal, generalized, and symplectic Stiefel manifolds mentioned above, and (3) How does an optimization method such as Riemannian gradient descent behave using the geometric quantities under the proposed metric?
+On the other hand, they inspire several questions: (1) How are we on track to avoid solving the Lyapunov matrix equations to compute the Riemannian gradient?
+(2) How to explain the different situations for orthogonal, generalized, and symplectic Stiefel manifolds mentioned above?
+And (3) how does an optimization method such as Riemannian gradient descent behave using the geometric quantities under the proposed metric?
 Contribution
@@
     $Z=XW+A^{-1}X_{\perp}K,$ (4)
-    where $W=JX^{T}AZ,K=(X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}Z,$ and $A^{-1}X_{\perp}K$ is independent of $X_{\perp}.$
+    where $W=JX^{T}AZ,K=(X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}Z$. It is immediately implied from (4) that $A^{-1}X_{\perp}K=Z-XW$ is independent of $X_{\perp}.$
   • iii)
@@
 3), a so-called tractable metric $\mathrm{g}_{\mathbf{M}_{X}}$ defined via an $X$-smoothly dependent matrix $\mathbf{M}_{X}\in\mathrm{SPD}(n)$ was proposed.
-Namely, for $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}(k,n),$ $\langle Z_{1},Z_{2}\rangle_{\mathrm{g}_{\mathbf{M}_{X}}}\coloneq\mathrm{tr}\left(Z_{1}^{T}\mathbf{M}_{X}Z_{2}\right)$.
+Namely, for any pair $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}(k,n),$ $\langle Z_{1},Z_{2}\rangle_{\mathrm{g}_{\mathbf{M}_{X}}}\coloneq\mathrm{tr}\left(Z_{1}^{T}\mathbf{M}_{X}Z_{2}\right)$.
 With this metric, for any $X\in\mathrm{iSt}(k,n)$, the ambient space $\mathbb{R}^{n\times k}$} is decomposed as the direct sum of the tangent space and the normal space,
@@
 $2\mathrm{sym}(X^{T}AY)=(J\Gamma_{1}J)U_{X,Y}+U_{X,Y}(J\Gamma_{1}J).$ (11)
-For a greater detail, let us convert this matrix equation into a linear equation using vectorization operator and Kronecker product HornJ91 .
-Namely, for a matrix $M=(m_{ij})\in{\mathbb{R}^{n\times k}}$, $\mathrm{vec}(M)\in\mathbb{R}^{nk}$ is generated by stacking the columns of $M$ to make a column vector and the Kronecker product of $M$ and $N\in\mathbb{R}^{p\times q}$, denoted by $M\otimes N\in\mathbb{R}^{np\times kq}$, is the matrix consisting of $nk$ block matrices $m_{ij}N$ for $i=1,\ldots,n,j=1,\ldots,k$.
+For greater detail, let us convert this matrix equation into a linear equation using vectorization operator and Kronecker product HornJ91 .
+Namely, for a matrix $M=(m_{ij})\in{\mathbb{R}^{n\times k}}$, $\mathrm{vec}(M)\in\mathbb{R}^{nk}$ is generated by stacking the columns of $M$ to make a column vector, and the Kronecker product of $M$ and $N\in\mathbb{R}^{p\times q}$, denoted by $M\otimes N\in\mathbb{R}^{np\times kq}$, is the matrix consisting of $nk$ block matrices $m_{ij}N$ for $i=1,\ldots,n,j=1,\ldots,k$.
 In view of (HornJ91, , Lem.
@@
 Thus, its solution is merely the product of the reciprocal of the scalar and the right-hand side.
-Accordingly, a sufficient condition for obviating the need of solving the Lyapunov matrix equation (8) is that $I_{k}\otimes(J\Gamma_{1}J)+(J\Gamma_{1}J)\otimes I_{k}$ has the form $2\rho I_{k^{2}}$, where $\rho$ is a nonzero constant parameter.
+Accordingly, a sufficient condition for obviating the need to solve the Lyapunov matrix equation (8) is that $I_{k}\otimes(J\Gamma_{1}J)+(J\Gamma_{1}J)\otimes I_{k}$ has the form $2\rho I_{k^{2}}$, where $\rho$ is a nonzero constant parameter.
 It follows that $J\Gamma_{1}J=\rho I_{k}$ which, together with the condition $J^{2}=I_{k}$, yields $\Gamma_{1}=\rho I_{k}$.
@@
 $\mathbf{M}_{X}^{-1}=E\begin{bmatrix}\rho I_{k}&0\\ 0&\Gamma_{3}\end{bmatrix}E^{T}=\rho XX^{T}+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}.$ (12)
-Moreover, taking (3) into account, we also have
+Moreover, taking the formula (3) for the inverse of $E$ into account, we also have
 $\mathbf{M}_{X}$ $=\begin{bmatrix}JX^{T}A\\ (X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}\end{bmatrix}^{T}\begin{bmatrix}\frac{1}{\rho}I_{k}&0\\ 0&\Gamma_{3}^{-1}\end{bmatrix}\begin{bmatrix}JX^{T}A\\ (X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}\end{bmatrix}$
@@
 Obviously, $g_{\rho,X_{\perp}}$ is a non-degenerate bilinear form defined on $T_{X}\mathrm{iSt}(k,n)$ and dependent upon both $\rho$ and $X_{\perp}$.
-To make this a Riemannian metric, except for the special parameter $\rho$, it is essential to require that $\mathbf{M}_{X}$ depends only on $X\in\mathrm{iSt}(k,n)$ and this dependence is smooth.
-To this end, we will impose additional conditions on $X_{\perp}$ in terms of $\Gamma_{3}$.
-If we choose
+To make this a Riemannian metric, it is essential to require that $\mathbf{M}_{X}$ depends only on $X\in\mathrm{iSt}(k,n)$ and this dependence is smooth.
+Apparently, we can consider $\Gamma_{3}$ as an independent parameter like $\rho$ and smoothly represent $X_{\perp}$ in terms of $X$ (or do this for both $\Gamma_{3}$ and $X_{\perp}$).
+Nevertheless, we do not know an efficient and explicit formula for this representation, if it exists.
+Following EdelAS98 ; GSAS21 , we impose a relation on these two components and obtain the desired metric.
+Namely, we can choose
 $\left(X_{\perp}^{T}X_{\perp}\right)^{-1}=\Gamma_{3}$ (15)
@@
 Remark 1
+  • The relations (15) and (17) are only the means to obtain $\mathbf{M}_{X}$. As we will see later, we do not have to construct $\Gamma_{3}$ or $X_{\perp}$.
   • One can see that when $\rho=2$ and $\Gamma_{3}=I_{n-k}$, the formulation (16) reduces to the canonical metric given in (EdelAS98, , Eq.
 (2.39)).
-Hence, the name generalized canonical metric.
+Hence, it is named the generalized canonical metric.
 Moreover, we use this value for $\rho$ in numerical tests unless stated otherwise.
@@
 3.1 Geometry under the generalized canonical metric
-For any $X\in\mathrm{iSt}(k,n),$ the matrix $\mathbf{M}_{X}$ defined by (13) is symmetric positive-definite, so we can extend the inner product $g_{\rho,X_{\perp}}$ on $T_{X}\mathrm{iSt}(k,n)$ to an inner product on $\mathbb{R}^{n\times k}$} which is given by $g_{\rho,X_{\perp}}(Z_{1},Z_{2})=\mathrm{tr}(Z_{1}^{T}\mathbf{M}_{X}Z_{2}),$ for all $Z_{1},Z_{2}\in{\mathbb{R}^{n\times k}}.$ The normal space to $\mathrm{iSt}(k,n)$ at $X$ with respect to $g_{\rho,X_{\perp}}$ is defined by
+For any $X\in\mathrm{iSt}(k,n),$ the matrix $\mathbf{M}_{X}$ defined by (13) is symmetric positive-definite, so we can extend the inner product $g_{\rho,X_{\perp}}$ on $T_{X}\mathrm{iSt}(k,n)$ to an inner product on $\mathbb{R}^{n\times k}$} which is given by $g_{\rho,X_{\perp}}(Z_{1},Z_{2})=\mathrm{tr}(Z_{1}^{T}\mathbf{M}_{X}Z_{2}),$ for all $Z_{1},Z_{2}\in{\mathbb{R}^{n\times k}}.$ For the moment, neither (15) nor (17) is imposed for the sake of generality.
+The normal space to $\mathrm{iSt}(k,n)$ at $X$ with respect to $g_{\rho,X_{\perp}}$ is defined by
 $T_{X}\mathrm{iSt}(k,n)_{\rho}^{\perp}\coloneqq\left\{N\in{\mathbb{R}^{n\times k}}:g_{\rho,X_{\perp}}(N,Z)=0\text{ for all }Z\in T_{X}\mathrm{iSt}(k,n)\right\}.$
@@
 $\mathrm{grad}_{gc}f(X)=\rho\,XJ\mathrm{skew}(JX^{T}\nabla\bar{f}(X))+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}\nabla\bar{f}(X).$ (22)
-Proof In view of the definition of the Riemannian gradient and (14), there holds
+Proof First, recall that the metric $g_{\rho,X_{\perp}}$ in (14) can be extended to $\mathbb{R}^{n\times k}$, and denote by $\mathrm{grad}_{gc}\bar{f}(X)$ the Riemannian gradient of $\bar{f}$ on $\mathbb{R}^{n\times k}$ equipped with that extended metric.
+In view of the definition of the Riemannian gradient, there holds
 $\langle\mathbf{M}_{X}\mathrm{grad}_{gc}\bar{f}(X),Z\rangle=g_{\rho,X_{\perp}}\left(\mathrm{grad}_{gc}\bar{f}(X),Z\right)=\textsc{D}\bar{f}(X)[Z]=\langle\nabla\bar{f}(X),Z\rangle,$
-for any $Z\in{\mathbb{R}^{n\times k}}.$ Therefore, from (12),
+for any $Z\in{\mathbb{R}^{n\times k}}.$ Therefore, from the formula (12) for the inverse of $\mathbf{M}_{X}$,
 $\mathrm{grad}_{gc}\bar{f}(X)=\mathbf{M}_{X}^{-1}\nabla\bar{f}(X)=\left(\rho XX^{T}+A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}\right)\nabla\bar{f}(X).$
@@
 $+A^{-1}\left(I_{n}-X(X^{T}X)^{-1}X^{T}\right)A^{-1}\nabla\bar{f}(X).$
-In the case that $\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}=\Gamma_{3}$, taking (5) into account, it follows that
+In the case that $\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}=\Gamma_{3}$, taking the formula (5) resulting from the identity $I_{n}=EE^{-1}$ into account, it follows that
 $A^{-1}X_{\perp}\Gamma_{3}X_{\perp}^{T}A^{-1}$ $=A^{-1}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}A^{-1}$
@@
 $+\left(I_{n}-XJX^{T}A\right)\left(I_{n}-XJX^{T}A\right)^{T}\nabla\bar{f}(X).$
-If matrix $A$ is dense and has no special structure, the computational cost for $A^{-1}(\cdot)$ in (23) is generally comparable to $n^{3}$.
-Therefore in this case, using (24) is more advantageous.
+If matrix $A$ is dense and has no additional special structure, the computational cost for $A^{-1}(\cdot)$ in (23) is generally comparable to $n^{3}$.
+Therefore, in this case, using (24) is more advantageous.
 One can see that the computation of (24) requires $2n^{2}k+10nk^{2}-nk-3k^{2}$ flops.
-Obviously, when $k$ is approaching $n$, this cost is cheaper than that caused by (9).
+Obviously, when $k$ is of the same scale as $n$, i.e., $k=\mathcal{O}(n)$, this cost is cheaper than that caused by (9).
 This will be seen in the numerical examples presented in Section 6.
@@
 (2.4)) via a change of variable.
-This explanation is also applicable to the geometry of the symplectic Stiefel manifold under the canonical-like metric GSAS21 as a special case of the tractable metric GSS2024_2 .
+In spite of the difference, similar arguments also work for the geometry of the symplectic Stiefel manifold under the canonical-like metric GSAS21 as a special case of the tractable metric GSS2024_2 .
 4 Quasi-geodesics with respect to the generalized canonical metric
@@
 In the dawn of Riemannian optimization, the solution is updated by searching along the geodesic which goes through the current solution in an appropriate tangent direction EdelAS98 ; Gaba82 ; Smit94 .
-In the Euclidean space, zero acceleration means $\ddot{X}(t)=0$ which implies that $X(t)=\dot{X}(0)t+X(0)$ which is a straight line.
+In the Euclidean space, zero acceleration means $\ddot{X}(t)=0$, or equivalently $X(t)=\dot{X}(0)t+X(0)$, which yields a straight line.
 That is, geodesics can be viewed as a generalization of lines in the Euclidean space to the Riemannian manifold.
@@
 Fortunately, in optimization, only local information encoded by geodesics and its feasibility are interesting due to their usage in updating the iterate.
-Thus, one usually retains these features in approximating geodesics which results in the so-called quasi-geodesics, see, e.g., JurdML2020 ; KrakMLB2017 ; NishA2005 for more approaches to the quasi-geodesics on the Stiefel manifold.
+Thus, one usually retains these features in approximating geodesics which results in the so-called quasi-geodesics; see, e.g., JurdML2020 ; KrakMLB2017 ; NishA2005 for more approaches to the quasi-geodesics on the orthogonal Stiefel manifold.
 In this paper, we follow a procedure similar to that of (GSAS21, , Subsect.
@@
 2.2).
-Namely, we replace the acceleration that should be formulated in accordance with the Riemannian connection associated with the proposed generalized canonical metric by the one corresponding to the Euclidean metric, which turns out to be $\ddot{Y}(t).$ In view of the normal vectors formulated in Proposition 1, we define the quasi-geodesic with respect to the generalized canonical metric $g_{\rho,X_{\perp}}$ to be any solution $Y_{\rho}(t)\in\mathrm{iSt}(k,n)$ of the initial value problem
+Namely, we replace the acceleration that should be formulated in accordance with the Riemannian connection associated with the proposed generalized canonical metric by the one corresponding to the Euclidean metric, which turns out to be $\ddot{Y}(t).$ In view of the normal vectors formulated in Proposition 1, we define a quasi-geodesic with respect to the generalized canonical metric $g_{\rho,X_{\perp}}$ to be any solution $Y_{\rho}(t)\in\mathrm{iSt}(k,n)$ of the initial value problem
 $\ddot{Y}(t)+Y(t)U(t)=0,\ Y(0)=X\in\mathrm{iSt}(k,n),\ \dot{Y}(0)=Z\in T_{X}\mathrm{iSt}(k,n),$ (25)
-with some $U(t)$ such that $JU\in\mathrm{Sym}(k)$ for all $t\geq 0$, although the obtained result apparently holds for $t<0$ too.
+with some $U(t)$ such that $JU(t)\in\mathrm{Sym}(k)$ for all $t\geq 0$, although the obtained result apparently holds for $t<0$ too.
 In this section, a closed-form solution to this equation is presented.
-For the sake of brevity, we refrain from writing the variable $t$ in our presentation.
+For the sake of brevity, we refrain from writing the variable $t$ in our presentation except for the arguments in which the dependence on $t$ is highlighted.
 As the first step, we show that for the equation (25) to have a solution, $U(t)$ must have a special form.
@@
 respectively. Substituting $\ddot{Y}=-YU$ from (25) into (27) gives
-$(-YU)^{T}AY+2\dot{Y}^{T}A\dot{Y}+Y^{T}A(-YU)=0$ (28)
-or $\dot{Y}^{T}A\dot{Y}=JU$ due to the symmetry of $JU$. Hence, $U(t)=J\dot{Y}^{T}A\dot{Y}$ and therefore, the equation (25) becomes
+$(-YU)^{T}AY+2\dot{Y}^{T}A\dot{Y}+Y^{T}A(-YU)=0$
+or $\dot{Y}^{T}A\dot{Y}=JU$ due to the symmetry of $JU$. Hence, $U=J\dot{Y}^{T}A\dot{Y}$ and therefore, the equation (25) becomes
 $\ddot{Y}+YJ(\dot{Y}^{T}A\dot{Y})=0.$ (28)
-Next, let us set $W\coloneqq Y^{T}A\dot{Y}$ and $V\coloneqq\dot{Y}^{T}A\dot{Y}$. Thanks to the relation (26), $W\in\mathrm{Skew}(k)$ for $t\geq 0.$ Moreover, using (29), we get
-$\dot{W}=V+Y^{T}A\ddot{Y}=V-Y^{T}AYJV=V-V=0,\mbox{ for }t\geq 0.$
-Hence, $W(t)=W(0)\eqqcolon W_{0}\in\mathrm{Skew}(k)$ for $t\geq 0$. We proceed to determine $V$ which is symmetric. Taking the derivative of $V$ with (29) in mind, we have that
+Next, let us set $S\coloneqq Y^{T}A\dot{Y}$ and $V\coloneqq\dot{Y}^{T}A\dot{Y}$. Thanks to the relation (26), $S\in\mathrm{Skew}(k)$ for $t\geq 0.$ Moreover, using (28), we get
+$\dot{S}=V+Y^{T}A\ddot{Y}=V-Y^{T}AYJV=V-V=0,\mbox{ for }t\geq 0.$
+Hence, $S(t)=S(0)\eqqcolon S_{0}\in\mathrm{Skew}(k)$ for $t\geq 0$. We proceed to determine $V$ which is symmetric. Taking the derivative of $V$ with (28) in mind, we have that
 $\dot{V}$ $=\ddot{Y}^{T}A\dot{Y}+\dot{Y}^{T}A\ddot{Y}=(-YJV)^{T}A\dot{Y}+\dot{Y}^{T}A(-YJV)$
-$=-VJW+WJV=W_{0}JV-VJW_{0}.$
+$=-VJS+SJV=S_{0}JV-VJS_{0}.$
 We will show that the initial value problem
-$\dot{V}=W_{0}JV-VJW_{0},\ V(0)=V_{0},$ (30)
-admits a solution $V(t)=\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}.$ Indeed, differentiating such $V(t)$ yields
-$\dot{V}$ $=-\mathrm{e}^{-tW_{0}^{T}J}W_{0}^{T}JV_{0}\mathrm{e}^{-tJW_{0}}-\mathrm{e}^{-tW_{0}^{T}J}V_{0}JW_{0}\mathrm{e}^{-tJW_{0}}$
-$=-W_{0}^{T}J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}-\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}JW_{0}$
-$=-W_{0}^{T}JV-VJW_{0}=W_{0}JV-VJW_{0}.$
-Here, we invoke the skew-symmetry of $W_{0}$ and the identity $M\mathrm{e}^{tM}=\mathrm{e}^{tM}M$ for any real square matrix $M$ and any real number $t$.
-Next, we present a technical lemma, which is similar to (GSAS21, , Lem. 4.6).
+$\dot{V}=S_{0}JV-VJS_{0},\ V(0)=V_{0},$ (29)
+admits a solution $V(t)=\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}.$ Indeed, differentiating such $V(t)$ yields
+$\dot{V}$ $=-S_{0}^{T}J\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}-\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}JS_{0}$
+$=-S_{0}^{T}JV-VJS_{0}=S_{0}JV-VJS_{0}.$
+Next, we present a technical lemma, which is similar to (GSAS21, , Prop. 4.6).
 Lemma 1
@@
 has a unique global solution in $\mathrm{iSt}(k,n)$ given by
-$Y_{\rho}(t)=Y_{\rho}^{\mathrm{qgeo}}(t,Z)\coloneqq[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$ (33)
-where $W_{0}=Y_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=X^{T}AZ$ and $V_{0}\coloneq\dot{Y}_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=Z^{T}AZ.$
+$Y_{\rho}(t)=Y_{\rho}^{\mathrm{qgeo}}(t,Z)\coloneqq[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}},$ (32)
+where $S_{0}=Y_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=X^{T}AZ$ and $V_{0}\coloneq\dot{Y}_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=Z^{T}AZ.$
 Proof First, let us assume that the problem (31) has a global solution in $\mathrm{iSt}(k,n),$ denoted also by $Y_{\rho}(t)$. It follows that
-$\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\right)$ $=\dot{Y_{\rho}}(t)\mathrm{e}^{tJW_{0}}+Y_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}.$
-$\frac{\mathrm{d}}{\mathrm{d}t}\left(\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right)$ $=\ddot{Y_{\rho}}(t)\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
-$=\left(-Y_{\rho}(t)JV\right)\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
-$=-Y_{\rho}(t)J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
-$=-Y_{\rho}(t)\mathrm{e}^{tJW_{0}}JV_{0}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0},$
-in which we have invoked (30) to replace $V\coloneq\dot{Y_{\rho}}^{T}A\dot{Y_{\rho}}$ by its explicit form $V=\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}$ when $Y_{\rho}(t)\in\mathrm{iSt}(k,n),$ for all $t\geq 0.$ Therefore,
-$\frac{\mathrm{d}}{\mathrm{d}t}\left[Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right]=\left[Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right]\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}$.
+$\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}(t)\mathrm{e}^{tJS_{0}}\right)$ $=\dot{Y_{\rho}}(t)\mathrm{e}^{tJS_{0}}+Y_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0}.$
+$\frac{\mathrm{d}}{\mathrm{d}t}\left(\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}\right)$ $=\ddot{Y_{\rho}}(t)\mathrm{e}^{tJS_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0}$
+$=\left(-Y_{\rho}(t)JV\right)\mathrm{e}^{tJS_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0}$
+$=-Y_{\rho}(t)J\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}\mathrm{e}^{tJS_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0}$
+$=-Y_{\rho}(t)\mathrm{e}^{tJS_{0}}JV_{0}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0},$
+in which we have invoked (29) to replace $V\coloneq\dot{Y_{\rho}}^{T}A\dot{Y_{\rho}}$ by its explicit form $V=\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}$ when $Y_{\rho}(t)\in\mathrm{iSt}(k,n),$ for all $t\geq 0.$ Therefore,
+$\frac{\mathrm{d}}{\mathrm{d}t}\left[Y_{\rho}(t)\mathrm{e}^{tJS_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}\right]=\left[Y_{\rho}(t)\mathrm{e}^{tJS_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}\right]\begin{bmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{bmatrix}$.
 (33)
 This basic equation admits an explicit solution from which we obtain that
-$Y_{\rho}(t)$ $=\left[Y_{\rho}(0)\;\;\dot{Y}_{\rho}(0)\right]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
-$=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$
+$Y_{\rho}(t)$ $=\left[Y_{\rho}(0)\;\;\dot{Y}_{\rho}(0)\right]\mathrm{exp}\left(t\begin{bmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
+$=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}},$
 which is the expected formula.
-Next, we will show that this is indeed a global solution to the problem (32).
-For brevity, we set $\Psi\coloneq\left[\begin{smallmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{smallmatrix}\right]$.
+Next, we will show that this is indeed a global solution in $\mathrm{iSt}(k,n)$ to the problem (31).
+For future reference, we present the feasibility part in Proposition 5 below.
+In the interest of brevity, we set $\Psi\coloneq\left[\begin{smallmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{smallmatrix}\right]$.
 In view of (32), it holds that
-$\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}+Y_{\rho}\mathrm{e}^{tJW_{0}}JW_{0}=\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}\mathrm{e}^{tJW_{0}}\right)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}$.
+$\dot{Y}_{\rho}\mathrm{e}^{tJS_{0}}+Y_{\rho}\mathrm{e}^{tJS_{0}}JS_{0}=\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}\mathrm{e}^{tJS_{0}}\right)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}$.
 This leads to
-$\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}-[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JW_{0}$ (35)
-$=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\biggr(\begin{bmatrix}JW_{0}\\ I_{k}\end{bmatrix}-\begin{bmatrix}JW_{0}\\ 0\end{bmatrix}\biggr)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
-Repeating this procedure with $\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}$, we obtain
-$\ddot{Y}_{\rho}\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}JW_{0}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
+$\dot{Y}_{\rho}\mathrm{e}^{tJS_{0}}=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}-[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JS_{0}$ (34)
+$=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\biggr(\begin{bmatrix}JS_{0}\\ I_{k}\end{bmatrix}-\begin{bmatrix}JS_{0}\\ 0\end{bmatrix}\biggr)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
+Repeating this procedure with $\dot{Y}_{\rho}\mathrm{e}^{tJS_{0}}$, we obtain
+$\ddot{Y}_{\rho}\mathrm{e}^{tJS_{0}}+\dot{Y}_{\rho}\mathrm{e}^{tJS_{0}}JS_{0}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
 Using (34), this identity gives
-$\ddot{Y}_{\rho}\mathrm{e}^{tJW_{0}}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}-JV_{0}\\ 0\end{bmatrix}$. (36)
+$\ddot{Y}_{\rho}\mathrm{e}^{tJS_{0}}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}-JV_{0}\\ 0\end{bmatrix}$. (35)
 On the other hand, using the expression for $\dot{Y}_{\rho}$ in (34), we receive the identities
-$\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho}=$ $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
-$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
-$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
-$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
-$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
-$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}},$ (37)
+$\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho}=$ $\left(\mathrm{e}^{-tJS_{0}}\right)^{T}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
+$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}J&S_{0}\\ -S_{0}&V_{0}\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
+$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
+$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
+$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}[0\;\;I_{k}]\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
+$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}},$ (36)
 in which Lemma 1 have been invoked in the fifth identity.
-Finally, left multiplying (37) by $Y_{\rho}J$ yields
-$Y_{\rho}J(\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho})=$ $\,Y_{\rho}J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}$
-$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}J^{2}\mathrm{e}^{tJW_{0}}JV_{0}\mathrm{e}^{-tJW_{0}}$
-$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JV_{0}\mathrm{e}^{-tJW_{0}}$
-$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}JV_{0}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$ (38)
+At the end, left multiplying (36) by $Y_{\rho}J$ yields
+$Y_{\rho}J(\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho})=$ $\,Y_{\rho}J\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}$
+$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}}J^{2}\mathrm{e}^{tJS_{0}}JV_{0}\mathrm{e}^{-tJS_{0}}$
+$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JV_{0}\mathrm{e}^{-tJS_{0}}$
+$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}JV_{0}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}},$ (37)
 in which the second identity is due to the fact that
-$(\mathrm{e}^{-tJW_{0}})^{T}=\mathrm{e}^{-tW_{0}^{T}J}=J^{2}\mathrm{e}^{-tW_{0}^{T}J}=J\mathrm{e}^{tJW_{0}}J.$
+$(\mathrm{e}^{-tJS_{0}})^{T}=\mathrm{e}^{-tS_{0}^{T}J}=J^{2}\mathrm{e}^{-tS_{0}^{T}J}=J\mathrm{e}^{tJS_{0}}J.$ (38)
 The conclusion follows from the two identities (35) and (37).
-Finally, the uniqueness of the solution is derived from that of the problems (30) and (34) which is a well-known fact in the theory of linear differential equations in Banach spaces, see, e.g., (Deimling1977, , Sect.
+The uniqueness of the solution is derived from that of the problems (29) and (33) which is a well-known fact in the theory of linear differential equations in Banach spaces; see, e.g., (Deimling1977, , Sect.
 1.1).
-∎
-The second part of the proof for Proposition 4, implicitly assumes that the curve $Y_{\rho}(t)$ in (33) is feasible. The statement below confirms it.
 Proposition 5
@@
 Proof From (32), it follows that
-$Y_{\rho}^{T}AY_{\rho}=\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[I_{k}\;\;0]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}.$ (39)
+$Y_{\rho}^{T}AY_{\rho}=\left(\mathrm{e}^{-tJS_{0}}\right)^{T}[I_{k}\;\;0]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}}.$ (39)
 According to the proof for (36), we have that
-$\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)=\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi=\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}$.
+$\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)=\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi=\begin{bmatrix}J&S_{0}\\ -S_{0}&V_{0}\end{bmatrix}$.
 This fact leads (39) to
-$Y_{\rho}^{T}AY_{\rho}=$ $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[I_{k}\;\;0]\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
-$=$ $J\mathrm{e}^{tJW_{0}}J^{2}\mathrm{e}^{-tJW_{0}}=J,$
-in which the second identity follows from the fact that $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}=J\mathrm{e}^{tJW_{0}}J.$
+$Y_{\rho}^{T}AY_{\rho}=$ $\left(\mathrm{e}^{-tJS_{0}}\right)^{T}[I_{k}\;\;0]\begin{bmatrix}J&S_{0}\\ -S_{0}&V_{0}\end{bmatrix}\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
+$=$ $J\mathrm{e}^{tJS_{0}}J^{2}\mathrm{e}^{-tJS_{0}}=J,$
+in which the second identity follows from the fact that $\left(\mathrm{e}^{-tJS_{0}}\right)^{T}=J\mathrm{e}^{tJS_{0}}J$ derived in (38). ∎
 In the case of generalized Stiefel manifold, i.e., $A$ is spd and $J=I_{k}$, the obtained quasi-geodesic reduces to a simple form as follows.
 Corollary 1
-Given $X\in\mathrm{St}_{A}(k,n)$, $Z\in T_{X}\mathrm{St}_{A}(k,n)$ and set $W_{0}=X^{T}AZ$, $V_{0}=Z^{T}AZ$. The curve
-$Y(t)=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}W_{0}&-V_{0}\\ I_{k}&W_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tW_{0}}$ (40)
+Given $X\in\mathrm{St}_{A}(k,n)$, $Z\in T_{X}\mathrm{St}_{A}(k,n)$, and set $S_{0}=X^{T}AZ$, $V_{0}=Z^{T}AZ$. The curve
+$Y(t)=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}S_{0}&-V_{0}\\ I_{k}&S_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tS_{0}}$ (40)
 is a quasi-geodesic associated with $X$ and $Z$ in the sense that $Y(t)\in\mathrm{St}_{A}(k,n)$ for all $t\geq 0$ and $Y(0)=X,\dot{Y}(0)=Z$.
@@
 2.2).
-It is worth to note that the one in (EdelAS98, , Subsect.
+It is worth noting that the one in (EdelAS98, , Subsect.
 2.2) is indeed associated with the Euclidean metric.
@@
 5 Quasi-geodesic based retraction on $\mathrm{iSt}(k,n)$
-Let us first formally define a retraction on $\mathrm{iSt}(k,n)$ AbsiMS08 ; AdleDMMS02 .
+Let us first formally define a retraction on $\mathrm{iSt}(k,n)$; see AbsiMS08 ; AdleDMMS02 for more details about retractions on a general Riemannian manifold.
 It is a smooth mapping $\mathcal{R}$ from the tangent bundle $T\mathrm{iSt}(k,n)\coloneqq\bigcup_{X\in\mathrm{iSt}(k,n)}T_{X}\mathrm{iSt}(k,n)$ to $\mathrm{iSt}(k,n)$ satisfying the following conditions for all $X\in\mathrm{iSt}(k,n):$
@@
 In TiepSon2025 , a retraction based on the Cayley transformation was proposed.
-In this work, based on the formulation of a quasi-geodesic (33), we propose another retraction which will be termed quasi-geodesic retraction as follows:
+In this work, based on the formulation of a quasi-geodesic (32), we propose another retraction, which will be termed quasi-geodesic retraction, as follows:
 $\mathcal{R}^{\mathrm{qgeo}}:T\mathrm{iSt}(k,n)$ $\longrightarrow\mathrm{iSt}(k,n)$
@@
 The map $\mathcal{R}^{\mathrm{qgeo}}$ given in (41) is a globally defined retraction on $\mathrm{iSt}(k,n).$
-Proof We only need to verify the conditions (R.1) and (R.2) since the expression in (41) is well-defined for any $X\in\mathrm{iSt}(k,n)$ and $Z\in T_{X}\mathrm{iSt}(k,n)$ thanks to Proposition 5.
+Proof We only need to verify the conditions (R.1) and (R.2) since the expression in (41) is well-defined for any $X\in\mathrm{iSt}(k,n)$ and $Z\in T_{X}\mathrm{iSt}(k,n)$ thanks to Proposition 5
 It holds that $\mathcal{R}_{X}^{\mathrm{qgeo}}(0)=[X\;\;0]\mathrm{exp}\left(\left[\begin{smallmatrix}0&0\\ I_{k}&0\end{smallmatrix}\right]\right)\left[\begin{smallmatrix}I_{k}\\ 0\end{smallmatrix}\right]\mathrm{e}^{0}=[X\;\;0]\left[\begin{smallmatrix}I_{k}\\ I_{k}\end{smallmatrix}\right]=X$.
@@
 $=Z,$
-where we have invoked the identity $\frac{\mathrm{d}t}{t}\mathrm{e}^{M(t)}|_{t=0}=\mathrm{e}^{M(0)}\dot{M}(0)$ for smooth one-parameter matrix $M(t)$ such that $M(0)\dot{M}(0)=\dot{M}(0)M(0)$, see (Wilcox1967, , Eq.
+where we have invoked the identity $\frac{\mathrm{d}t}{t}\mathrm{e}^{M(t)}|_{t=0}=\mathrm{e}^{M(0)}\dot{M}(0)$ for smooth one-parameter matrix $M(t)$ such that $M(0)\dot{M}(0)=\dot{M}(0)M(0)$; see (Wilcox1967, , Eq.
 (2.1)).
@@
 6 Numerical examples
-In this section, theoretical results are assembled to enable a Riemannian gradient descent scheme and then numerically verified via examples.
+In this section, several numerical examples are presented to verify our theoretical findings.
+As constructed in Section 3, two formulae for computing the Riemannian gradient using the proposed metric are given in (23) and (24); it was also stated that the latter outperforms the former when $A$ is dense.
+First, we validate this statement via an extra example.
+After that, the derived components are assembled to enable a Riemannian gradient descent scheme and then numerically verified via several model tests.
+The tests are done on a laptop with Intel(R) Core(TM) i7-4500U (at 1.8GHz, 5MB Cache, 8GB RAM) running MATLAB R2023a under Windows 10 Home.
+(The main code is made available at https://sites.google.com/view/ntson/code.)
+A numerical comparison of (23) and (24)
+We construct a manifold $\mathrm{iSt}_{A,J}$ as follows: the matrix $A=P\mathrm{diag}(1,2,\ldots,1500,-500,-499,\ldots,-1)P^{T}$, where $P$ is an orthogonalized random matrix and $J=\mathrm{diag}(I_{14},-I_{6})$.
+We take $f(X)=\frac{1}{2}\mathrm{tr}\left(X^{T}MX\right)$ with $M$ is a fixed spd matrix of apt size.
+We compute the Riemannian gradient using two formulae (23) and (24) at $X=\left[p_{1}/\sqrt{1},\ldots p_{14}/\sqrt{14},p_{1501}/\sqrt{500},\ldots,p_{1506}/\sqrt{495}\right]$ in which $p_{j}$ is the $j$-th column of $P$ generated previously.
+We run this setting 10 times and take the average time elapsed.
+The result shows that, given the Euclidean gradient $\nabla f(X)=MX$, while computing the Riemannian gradient following (23) takes about 0.45(s), the one invoking (24) needs only about 0.016(s), which is roughly 30 times faster.
+Therefore, only (24) is further used in our tests.
 A Riemannian gradient descent method
-To minimize a continuously differentiable cost function $f$ on the indefinite Stiefel manifold $\mathrm{iSt}_{A,J}$,
+One option to minimize a continuously differentiable cost function $f$ on the indefinite Stiefel manifold $\mathrm{iSt}_{A,J}$,
 $\min_{X\in\mathbb{R}^{n\times k}}f(X)\ \mbox{ s.t. }\ X^{T}AX=J,$ (42)
-given an initial point $X_{0}\in\mathrm{iSt}_{A,J}(k,n)$, one consecutively updates the approximate solution by
+given an initial point $X_{0}\in\mathrm{iSt}_{A,J}(k,n)$, is to consecutively updates the approximate solution by
 $X_{j+1}=\mathcal{R}_{X_{j}}(-\tau_{j}\mathrm{grad}f(X_{j})),$
 where $\mathcal{R}_{X_{j}}(\cdot)$ is a retraction at $X_{j}$, $\mathrm{grad}f(X_{j})$ denotes the Riemannian gradient of the cost function $f$ at $X_{j}$ with respect to a chosen Riemannian metric $\mathrm{g}(\cdot,\cdot)$, and $\tau_{j}>0$ is a step size to be determined.
-To compute the step size, we combine a nonmonotone line search technique with the alternating Barzilai-Borwein strategy on an embedded submanifold, see BarB88 ; HuLWY2020 ; IannP18 ; ZhanH2004 .
+To compute the step size, we combine a nonmonotone line search technique with the alternating Barzilai-Borwein strategy on an embedded submanifold; see BarB88 ; HuLWY2020 ; IannP18 ; ZhanH2004 .
 For the sake of convenience, we recall the details in Algorithm 1.
@@
 Moreover, $\alpha=0.85$ is used in the nonmonotone condition.
-A running is considered convergent when the norm of the Riemannian gradient of the cost function at the current iterate is relatively small compared to that at the initial guess $X_{0}$, namely, $\|\mathrm{grad}f(X_{j})\|_{X_{j}}\leq\texttt{rstop}\|\mathrm{grad}f(X_{0})\|_{X_{0}}$.
+A running is considered convergent when the norm of the Riemannian gradient of the cost function at the current iterate is relatively small compared to that at the initial guess $X_{0}$.
+Namely, $\|\mathrm{grad}f(X_{j})\|_{X_{j}}\leq\texttt{rstop}\|\mathrm{grad}f(X_{0})\|_{X_{0}}$, where $\|\cdot\|_{X_{j}}$ is the norm associated with the chosen metric at $X_{j}$.
 In all tests, rstop is set to be $1\mathrm{e}-5$.
+We also stop the loop when $j$ exceeds a given maximal number of iterations iter which is set to $1000$.
 Model problems
@@
 $f(X):=\mathrm{tr}(X^{T}MX),\ X\in\mathrm{iSt}_{A,J}(k,n),$ (44)
-where $n=1000,k=200$, $M=VV^{T}$ with $V$ is the $n\times(n-5)$ matrix randomly generated with default setting and then orthogonalized by orth in MATLAB.
-For the manifold, let $p=750,m=250,k_{p}=k_{m}=100$, we set $A=\mathrm{diag}(1,\ldots,p,-1,\ldots,-m)$ and $J=\mathrm{diag}(I_{k_{p}},-I_{k_{m}})$.
-The initial guess $X_{0}$ is set as the $n\times k$ zero matrix except for the submatrices $X_{0}(1,\ldots,k_{p};1,\ldots,k_{p})=\mathrm{diag}(\frac{1}{\sqrt{1}},\ldots,\frac{1}{\sqrt{k_{p}}})$ and $X_{0}(p+1,\ldots,p+k_{m};k_{p}+1,\ldots,k_{p}+k_{m})=\mathrm{diag}(\frac{1}{\sqrt{1}},\ldots,\frac{1}{\sqrt{k_{m}}})$.
+where $n=1000,k=500$, $M=VV^{T}$ with $V$ is the $n\times(n-5)$ matrix randomly generated with default setting and then orthogonalized using orth in MATLAB; matrix $M$ is finally normalized by dividing $M$ by its Frobenius norm.
+For the manifold, let $p=750,m=250,k_{p}=400,k_{m}=100$, we set $A=\mathrm{diag}(1,\ldots,p,-1,\ldots,-m)$ and $J=\mathrm{diag}(I_{k_{p}},-I_{k_{m}})$.
+The initial guess $X_{0}$ is, for the sake of simplicity, set as the $n\times k$ zero matrix except for the submatrices $X_{0}(1,\ldots,k_{p};1,\ldots,k_{p})=\mathrm{diag}(1/\sqrt{1},\ldots,1/\sqrt{k_{p}})$ and $X_{0}(p+1,\ldots,p+k_{m};k_{p}+1,\ldots,k_{p}+k_{m})=\mathrm{diag}(1/\sqrt{1},\ldots,1/\sqrt{k_{m}})$.
 The second model is the well-known Procrustes problem
 $f(X)=\|GX-B\|^{2}_{F}$ (45)
-on the $J$-orthogonal group with $n=500$, $J=\mathrm{diag}(I_{p},-I_{m}),$ where $p=375,$ $m=125$.
-Numerical data are generated as follow: $G$ is a $(n-20)\times n$ random matrix with default setting and $B$ is set as $G\mathrm{diag}(U,V)$ where $U\in\mathbb{R}^{p\times p}$ and $V\in\mathbb{R}^{m\times m}$ are two orthogonalized random matrices generated with the setting rng (1,’twister’).
+on the $J$-orthogonal group High03 with $n=500$, $J=\mathrm{diag}(I_{p},-I_{m}),$ where $p=375,$ $m=125$.
+Numerical data are generated as follows: $G$ is a $(n-20)\times n$ random matrix with default setting and then normalized as above, and $B$ is set as $G\mathrm{diag}(U,V)$ where $U\in\mathbb{R}^{p\times p}$ and $V\in\mathbb{R}^{m\times m}$ are two orthogonalized random matrices generated with the setting rng (1,’twister’).
 The $n\times n$ identity matrix is used as the initial guess.
 Tests
-It is worth to note that, while the Riemannian gradient is defined upon the metric, the retraction is metric-independent.
+It is worth noting that, while the Riemannian gradient is defined upon the metric, the retraction is metric-independent.
 Therefore, in Algorithm 1, one can combine any metric with any available retraction.
@@
 Regarding metrics, we have the general tractable and the generalized canonical metrics.
-In TiepSon2025 , the Euclidean and weighted Euclidean metrices have been used.
+In TiepSon2025 , the Euclidean and weighted Euclidean metrics have been used.
 The examples there showed that using the latter is much more advantageous.
@@
 Consequently, there are totally four combinations whose corresponding numerical results will be marked by “Eucl-Cayley”, “Eucl-qgeo”, “gcan-Cayley”, and “gcan-qgeo”.
-Besides, since in our code we do not attempt to exploit the special structure of $A$, we always employ (17) for $X_{\perp}$.
-All the tests are done on a laptop with Intel(R) Core(TM) i7-4500U (at 1.8GHz, 5MB Cache, 8GB RAM) running MATLAB R2023a under Windows 10 Home.
-(The main code is made available at https://sites.google.com/view/ntson/code.)
-We report the numerical results in Table 1 and Table 2 in which obj., grad., feas., #\#iter., #\#eval., and CPU represent the value of the cost function, the norm in the chosen metric of the Riemannian gradient of the cost function, the feasibility $\|X_{*}^{T}AX_{*}-J\|_{F}$, the number of iterations, the number of function evaluations, and the elapsed time, respectively.
+We report the numerical results in Table 1 and Table 2 in which obj., grad., feas., #\#iter., #\#eval., and CPU represent the value of the cost function, the norm in the chosen metric of the Riemannian gradient of the cost function, the feasibility $\|{X^{*}}^{T}AX^{*}-J\|_{F}$, the number of iterations, the number of function evaluations, and the elapsed time in second, respectively.
 In the case that the iteration has not reached the stopping criterion, $X^{*}$ is replaced by the last iterate.
-In addition, the convergence histories are displayed in Figure 1 and Figure 2.
-Table 1: Results for minimization of trace cost function (44).
+In addition, the convergence histories are displayed in Figure 1 and Figure 2 in which we present the value of the cost function, the norm of the Riemannian gradient, and the time elapsed along the iterations, respectively from the left to the right.
+Table 1: Results for minimization of the trace cost function (44).
 Combination obj. grad. feas. #\#iter. #\#eval. CPU
-Eucl-Cayley $0.6387$ $4.0204\mathrm{e}-6$ $2.3354\mathrm{e}-13$ 389 423 97.3941
-Eucl-qgeo $0.6387$ $3.9532\mathrm{e}-6$ $5.0332\mathrm{e}-08$ 392 437 50.3922
-gcan-Cayley 0.6387 $4.2671\mathrm{e}-6$ $2.5933\mathrm{e}-13$ 472 516 93.2189
-gcan-qgeo 0.6387 $4.1930\mathrm{e}-6$ $3.1692\mathrm{e}-13$ 458 598 47.1033
-Table 2: Results for minimization of Procrustes cost function (45).
+Eucl-Cayley $0.0398$ $1.0383\mathrm{e}-7$ $5.1411\mathrm{e}-13$ 480 524 750
+Eucl-qgeo $0.0398$ $1.2258\mathrm{e}-7$ $9.9156\mathrm{e}-08$ 407 446 685
+gcan-Cayley 0.0398 $1.0342\mathrm{e}-7$ $6.6291\mathrm{e}-13$ 614 666 260
+gcan-qgeo 0.0398 $1.1001\mathrm{e}-7$ $4.8909\mathrm{e}-13$ 461 505 324
+Table 2: Results for minimization of the Procrustes cost function (45).
 Combination obj. grad. feas. #\#iter. #\#eval. CPU
-Eucl-Cayley $21.5943$ $5.9039$ $2.5425\mathrm{e}-12$ 2000 2042 $2068$
-Eucl-qgeo $38.1905$ $5.6773$ $8.4230\mathrm{e}-10$ 2000 2064 $2817$
-gcan-Cayley 2.8671 $1.4755$ $3.3286\mathrm{e}-13$ 340 615 54
-gcan-qgeo 15.5090 $1.5283$ $1.2650\mathrm{e}-10$ 402 738 356
+Eucl-Cayley $9.0163\mathrm{e}-4$ $1.9212\mathrm{e}-3$ $1.0842\mathrm{e}-12$ 1000 1034 $1096$
+Eucl-qgeo $7.6519\mathrm{e}-4$ $2.0001\mathrm{e}-4$ $8.4243\mathrm{e}-13$ 1000 1034 $1491$
+gcan-Cayley $2.0092\mathrm{e}-4$ $2.0370\mathrm{e}-5$ $4.2354\mathrm{e}-13$ 435 757 70
+gcan-qgeo $1.9463\mathrm{e}-4$ $1.8871\mathrm{e}-5$ $6.4464\mathrm{e}-13$ 402 734 400
 Figure 1: Convergence history of the Riemannian gradient descent method with different combinations of metrics and retractions applied to the trace cost function (44).
@@
 Discussion
-Through examples presented, we can first see that, all newly established theoretical materials work properly and compatibly with the existing ones.
-Second, as expected, for the case that $k$ is not small, the schemes using the generalized canonical metric always need less time to be convergent (with respect to the given rstop) compared to the ones using the Euclidean metric; even in the case of Procrustes problem, within at most 2000 iterations, only the schemes with the generalized canonical metric converge.
+Through examples presented, we can first see that, all newly established theoretical components work properly and compatibly with the existing ones.
+Second, as expected, for the cases that $k$ is not small in both models, the schemes using the generalized canonical metric need less time to be convergent with respect to a given rstop compared to the ones using the Euclidean metric although in some cases the required number of iterations can be larger.
 Third, the quasi-geodesic retraction apparently is less feasible than the Cayley retraction.
 This fact might come from the instability of the scaling and squaring algorithm invoked by expm in MATLAB which occasionally occurs as pointed out in GuetN2016 .
+And finally, we note that while our sole purpose in constructing the generalized canonical metric is to avoid solving the Lyapunov equation in computing the Riemannian gradient, altering the metric has a great effect on the aspects of numerical methods, such as (numerical) convergence, number of iterations, as one can see in our tests.
+The result can also depend on the model, namely the cost function.
+It is advised in MishS2016 that the Hessian of the cost function should be taken into account when constructing the metric.
+To understand the details, we believe that deeper and problem-dependent investigations are required.
 7 Conclusions